### v1

# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

Abstract.
Computational offloading is a promising approach for overcoming resource constraints on client devices by moving some or all of an application’s computations to remote servers. With the advent of specialized hardware accelerators, client devices are now able to perform fast local processing of specific tasks, such as machine learning inference, reducing the need for offloading computations. However, edge servers with accelerators also offer faster processing for offloaded tasks than was previously possible. In this paper, we present an analytic and experimental comparison of on-device processing and edge offloading for a range of accelerator, network, and application workload scenarios, with the goal of understanding when to use local on-device processing and when to offload computations. We present models that leverage analytical queuing results to capture the effects of dynamic factors such as the performance gap between the device and edge server, network variability, server load, and multi-tenancy on the edge server. We experimentally demonstrate the accuracy of our models for a range of hardware and application scenarios and show that our models achieve a mean absolute percentage error of 2.2% compared to observed latencies. We use our models to develop an adaptive resource manager for intelligent offloading and show its efficacy in the presence of variable network conditions and dynamic multi-tenant edge settings.
submissionid: 696copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXX
1. Introduction
Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment. In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements. Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference (Mosqueira-Rey et al., 2022). Edge computing, which is a complementary approach to cloud computing, is a promising method for meeting the needs of such latency-sensitive applications. Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing. However, recent research (Ali-Eldin et al., 2021; Wang et al., 2024a) has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud. In particular, since edge clusters are resource-constrained relative to the cloud, research has shown resource bottlenecks can arise in the presence of time-varying workloads and increase the processing latency at the edge, negating its network latency advantage. Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing (Wang et al., 2024a).
A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge. [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper. Computational offloading (Satyanarayanan et al., 2009; Davis et al., 2004) involves using nearby edge resources to take on some, or all, of an application’s processing demands. Such offloading is performed to overcome resource constraints on a local device by leveraging more abundant processing capacity at edge servers. Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers (Ren et al., 2019). In this case, local devices pay a so-called mobility penalty (Satyanarayanan et al., 2019), which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead. Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks (Jang et al., 2018). In these offloading scenarios, conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought. For example, the rise of the so-called AI PC with onboard AI accelerators (Intel Corporation, 2025) has enabled local AI processing for many applications without a need for edge offloading. Similarly, mobile phones are equipped with neural accelerators (Simonite, 2017) that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally (Netatmo, 2025). At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before. Thus, in the era of programmable accelerators, the question of whether to offload computations from local devices to the edge and when it is advantageous to do so needs to be reexamined. In particular, such a reexamination needs to address several questions. (i) Under what scenarios will local on-device processing using accelerators outperform edge offloading? (ii) Are there still scenarios where edge offloading has an advantage over local processing? (iii) What role does a fast (or slow) network between the device and the edge play in terms of whether to offload or not offload?
Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators. Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and the choice of where to perform the processing will depend on hardware, network, and workload characteristics. In developing our analytic models and validating this hypothesis, this paper makes the following contributions.
  • We use queuing theory-based models to model the behavior of device and edge accelerators when processing requests for local on-device processing and edge offloading scenarios. Using these models, we develop closed-form analytic bounds to determine when one approach outperforms the other under different hardware, network, and workload scenarios. We also develop analytic models for comparing on-device processing to multi-tenant edge servers that service multiple clients.
  • We experimentally validate our analytic models and bounds using several different device and edge accelerators and network configurations and show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios. Our evaluation shows that our model achieves a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
  • We develop a dynamic resource manager that incorporates our models and demonstrate the ability of this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios. Our results show that it effectively adapts to changing network conditions and fluctuating request rates at edge servers.
2. Background
This section provides background on computational offloading, hardware accelerators, and on-device processing.
Table 1. Characteristics of common programmable accelerators.
Accelerator Peak Power (W) Memory Cost Target Workload
Intel Movidius Myriad X VPU 1 – 2.5 2.5 MB of SRAM \$59 Computer vision and CNN inference
Google Edge TPU 2 8 MB of SRAM \$75 ML Inference with TensorFlow Lite
NVIDIA Jetson TX2 15 8 GB of LPDDR4 \$399 AI and GPU workload
NVIDIA Jetson Orin Nano 15 8 GB of LPDDR5 \$499 AI and GPU workload
NVIDIA A2 GPU 60 16 GB of GDDR6 \$1399 AI and GPU workload
2.1. Cloud and Edge Offloading
Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server. In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks (Satyanarayanan et al., 2009; Davis et al., 2004). While such edge offloading involves a network hop to the edge server (also known as the mobility penalty (Satyanarayanan et al., 2019)), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices. In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload (Huang et al., 2020). For example, large language model (LLM) inference uses GPU clusters in the cloud to answer user queries, while IoT devices use cloud storage and processing to perform analytic and data processing tasks.
In general, computational offloading offers benefits when local processing is constrained and the cost of incurring the additional network latency outweighs the benefits of accessing faster processing in the edge or the cloud.
2.2. Hardware Accelerators
In recent years, a wide range of programmable accelerators has been developed to accelerate a range of compute tasks, particularly machine learning compute. Such accelerators have been deployed on client devices, edge, and cloud servers, enabling faster processing of accelerated tasks at each tier. Table 1 depicts the characteristics of several commonly deployed device and edge accelerators, ranging from low-end to high-end options. Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications and CNN inference (Intel, 2017). Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge (Google, 2019). NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision (NVIDIA, 2017, 2023, 2021). For these specific workloads, programmable accelerators can provide substantial performance improvements, delivering server-like capabilities for some compute-intensive tasks even on low-end devices. With their small form factor and energy-efficient design, integrating accelerators into mobile devices has become increasingly common. For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device (Simonite, 2017). With the presence of accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without the need to resort to edge offloading. At the same time, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks. This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
2.3. On-device Processing versus Edge Offloading
As noted above, the availability of accelerators on client devices has enabled on-device processing of many tasks without the need to offload those tasks to the edge or to the cloud. Since local processing does not incur any network hops, it will often be faster than remote processing which incurs an additional network latency. However, since edge servers can also be equipped with GPUs and other accelerators, in some compute-intensive scenarios, edge processing may still provide benefits over local processing on device accelerators. Thus, in the era of accelerators, local on-device processing may be faster in some scenarios, while offloading to edge accelerators may still be faster in other cases (after accounting for the network latency).
Even with extensive offline profiling of execution times on both the device and the edge, selecting the optimal execution strategy remains challenging due to dynamic real-world conditions. On the one hand, mobile devices are subject to network variability due to fluctuating wireless signal strength, interference, and congestion in shared communication channels. Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting the overall offloading performance. Additionally, mobile devices running applications with prolonged operating durations, such as live object detection on captured video, may experience battery constraints. When operating in low-power mode, the system may underclock its processing units to conserve energy, resulting in extended execution times. On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can result in interference across users and higher latencies. Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically based on the number and nature of tasks being offloaded at any given time. Compared to traditional cloud data centers, edge servers are often more resource-constrained, limiting their ability to elastically scale applications under resource pressure. Workload spikes in multi-tenant scenarios can increase queuing delays, significantly increasing end-to-end execution times. Moreover, edge servers are commonly multiplexed across different applications to maximize edge resources utilization. At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs (NVIDIA, 2025), leading to performance interference between applications. As a result, execution times can become variable, making it challenging to determine whether to offload under dynamic workloads.
The above observations motivate the primary research question addressed in this paper: Considering the hardware processing capabilities of the device and edge server, along with the workload characteristics and real-world dynamics, under what scenarios does one strategy outperform the other? Specifically, does edge offloading still have a role to play when devices are equipped with on-board accelerators? How do workload dynamics, network dynamics, and edge multi-tenancy impact each strategy and their relative performance with respect to one another? To address this question, the next section develops analytic models to capture the end-to-end execution time for each on-device and edge processing, capturing the key factors that impact application performance.
3. Model-driven Performance Comparison
In this section, we develop analytical queuing models for edge offloading and on-device processing and then develop closed-form expressions for their expected latencies to compare the two approaches and determine when one strategy outperforms the other.
3.1. Analytic Queuing Models
(a) Edge offloading queueing model
(b) On-device processing queueing model
Figure 1. Modeling request execution using (a) edge offloading and (b) on-device processing.
Our analytical models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems. Queuing theory has previously been used to model the performance of edge and cloud applications (Ali-Eldin et al., 2021; Harchol-Balter, 2013; Urgaonkar et al., 2005), and will serve as the foundation for analyzing edge offloading and on-device processing. As discussed below, we model the end-to-end latency of a request—the duration of time from when the task enters the first queue until it exits the system—for both edge offloading and on-device processing using queuing theory.
Edge offloading. To model the end-to-end request execution time using edge offloading, we need to model the times spent by the request both at the device and at the edge server. We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1(a). We assume that the request arrives at the device and is then forwarded (“offloaded”) to the edge server over the network. We need to model any network queuing delays and the network transmission time of the request from the device to the server. Upon arrival at the edge server, the request is queued for processing and is subsequently scheduled for execution. Since accelerators employ data parallelism to process requests, we model the accelerator at the edge server as a parallel system that processes each request with a degree of parallelism $k_{edge}$. Once the request processing completes, the result is sent back to the device, which incurs network queuing delay and transmission latency on the reverse path.
On-device processing. Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1(b). Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores of the device where they are processed with a degree of parallelism $k_{dev}$. Upon completion of processing, the result is returned to the local application and the request exits the system. This model captures an application where a client device generates tasks, utilizes its accelerator cores for local execution, and sends the result to the application.
In queuing theory, the processing time of a request at each queue that it encounters consists of two key components: queuing delay (i.e., the time spent waiting in the queue) and service time (i.e., the time taken to execute the request on the accelerator). Let $w_{dev}^{proc}$, $w_{edge}^{proc}$, $s_{dev}$, and $s_{edge}$ represent the wait times and request service times at the device and edge, respectively. Edge offloading additionally incurs network latency to transmit the request to the edge and the response back to the device, while on-device processing does not incur any network latency since all processing is local. Let $w_{dev}^{net}$ and $w_{edge}^{net}$ denote the respective network queuing delays on the device and the edge. Also, let $n_{req}$ denote the network delay of sending a request from the device to the edge, and $n_{res}$ be the network delay of sending a result from the edge to the device.
With the above notations, we now model the end-to-end latency of requests executed using each strategy. First, we have
(1) $T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
and
(2) $T_{dev}=w_{dev}^{proc}+s_{dev}$
where $T_{edge}$ and $T_{dev}$ denote the end-to-end latency of requests executed using edge offloading and on-device processing, respectively.
3.2. Deriving and Analyzing End-to-End Latency
In this section we compute the expected latency for edge offloading and on-device processing. Our analysis focuses on applications that require processing results on the device, such as AR/VR applications. Note that the proposed analytic models can also capture the performance of applications requiring results on the edge, and our findings can be extended to these cases. In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
3.2.1. Comparing Edge Offload and On-Device Execution
We begin our analysis by examining the conditions under which edge offloading results in higher latency compared to on-device processing. For workloads that benefit from hardware acceleration, such as those utilizing GPUs or TPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs. In the following discussion, we refer to such workloads as accelerator-driven workloads. Table 2 summarizes the notation used in our models.
Table 2. Notations used in analytic models.
Notation Description
$T_{dev}$ End-to-end latency of on-device processing
$T_{edge}$ End-to-end latency of edge offloading
$\lambda$ Request arrival rate
$B$ Bandwidth of the device-to-edge network path
$D_{req},D_{res}$ Payload sizes of request/result
$w_{dev}^{net},w_{edge}^{net}$ Network wait time at the device/edge
$n_{req},n_{res}$ Network transmission time of request/response
$\mu_{dev}^{net},\mu_{edge}^{net}$ Service rate of device/edge NIC
$k_{dev},k_{edge}$ Parallelism level of processors at device/edge
$w_{dev}^{proc},w_{edge}^{proc}$ Processing wait time at the device/edge
$s_{dev},s_{edge}$ Service time at the device/edge
$\mu^{proc}_{dev},\mu^{proc}_{edge}$ Service rate of device/edge
Lemma 3.0.
For accelerator-driven workloads, edge offloading incurs a larger average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
Proof.
Edge offloading will offer a worse latency than on-device processing when $T_{device}<T_{edge}$. That is
(3) $w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
Let $D_{req}$ and $D_{res}$ denote the task and result payload size, respectively. Further, let $B$ denote the bandwidth of the device-to-edge network path. The above inequality reduces to
(4) $s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$
As shown in (Liang et al., 2023), workloads on programmable accelerators typically exhibit deterministic service times (denoted as $D$), as the number of operations per request remains constant. Since typical accelerators have a small degree of parallelism $k$ (Amert et al., 2017), the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process. While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ (Shah et al., 2018, 2020). This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$. Its expected wait time can be derived from the well-known Pollaczek-Khinchin (P-K) formula (Gross et al., 2008; Harchol-Balter, 2013) and is given by
(5) $E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$
On the other hand, network interfaces with a single controller can be modeled as $M/M/1/FCFS$ systems with expected wait times given by
(6) $E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$
Let $k_{dev}$ and $k_{edge}$ denote the degree of parallelism seen at the accelerator on the device and the edge respectively. Applying these two results to (4), we get {dmath} s_dev - s_edge ¡ 1μdevnet- λdevnet - 1μdevnet + 1μedgenet- λedgenet - 1μedgenet + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λedge - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
Offloading is only required when a request arrives, which gives us $\lambda^{net}_{dev}=\lambda$. For a queuing system to be stable, throughput (i.e., task completion rate) must equal the task arrival rate $\lambda$ under steady-state conditions. When tasks are offloaded to an edge server, results are transmitted to the device only upon task completion, which yields $\lambda_{edge}^{net}=\lambda$. The first two terms involving $\mu_{dev}^{net}$ on the RHS can be written as {dmath} 1μdevnet- λdevnet - 1μdevnet = λμdevnet(μdevnet- λ) whereas the second two terms involving $\lambda_{edge}^{net}$ can be written as {dmath} 1μedgenet- λedgenet - 1μedgenet = λμedgenet(μedgenet- λ).
Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$. Substituting the above into (3.2.1), we obtain: {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof. ∎
The above result implies the following two remarks.
Remark 3.2.
On-device processing is likely to outperform edge offloading for workloads with lower computational demand.
Explanation. Service time, which is the time spent processing a request on an accelerator, is determined by the computational demand of the workload divided by the processing capacity of the hardware. For any given hardware configuration, requests with lighter demand have to be processed much faster at the edge to overcome the network penalty for edge offloading to be faster than on-device processing. For requests with lower computational demands, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally due to the disparity in processing capacities between the device and the edge as the workload computational demand decreases. Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (3.1) will also become smaller since $\mu$ increases for lighter requests. The terms representing the network overheads remain unchanged since they depend on request size rather than the processing demand of the request. As a result, the inequality in (3.1) is more likely to hold true for requests with lighter processing demand. ∎
Remark 3.3.
On-device processing is more likely to outperform edge offloading for slower networks or when workloads have larger request or result payloads.
Explanation. The term $\frac{D_{req}+D_{res}}{B}$ in (3.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality. In addition, longer transmission times reduce the service rates of the NIC at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the RHS. As a result, the inequality is more likely to hold for slower networks or tasks with higher network overheads, making on-device processing faster than edge offloading. ∎
Practical takeaways. Lemma 3.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective. The bound and the resulting inequality depends on several factors: (i) relative processing speeds of the device and edge, denoted by $s_{dev}$, $s_{edge}$ (and also the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) network speed, (iii) overall workload denoted by request rate $\lambda$, (iv) network transmission overheads. In general, for any specific workload, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading the workload; otherwise, on-device processing becomes preferable to offloading.
Moreover, the request rate $\lambda$ impacts offloading decisions in two competing ways. First, as $\lambda$ increases, both the edge and the device experience longer queuing delays, with the edge queue growing more slowly due to its faster processing speed. Consequently, the benefits of offloading improve under high load. However, under slow network conditions or for applications with large payloads, a high $\lambda$ increases the network queuing delay (first two terms of the right side of the inequality) and may lead to network congestion, making on-device processing preferable. Thus, the optimal strategy depends on the workload’s payload size relative to both network bandwidth and the edge-device performance gap.
Remark 3.2 shows how the computational demand of the workload influences the execution strategy. For workloads with higher computational demand, edge offloading is more likely to outperform on-device processing, as the edge’s processing speed advantage becomes larger. Conversely, for workloads with lower computational demand, on-device processing is more efficient as the performance advantage of edge processing is likely outweighed by the associated network costs. Remark 3.3 demonstrates how network delays affect the efficiency of edge offloading. With smaller network bandwidth or larger request/result payloads, both the network queuing delay and transmission time increase. As a result, edge offloading becomes less efficient, while on-device processing is unaffected since processing is local and does not rely on network resources.
3.3. Impact of Multi-Tenancy
Figure 2. $m$ devices offloading requests to an edge server.
While the analysis in the Sec 3.2.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work. Hence, we extend our model-driven comparison to multi-tenancy scenarios.
Modeling edge application multiplexing. In the case where each edge server is dedicated to a single client device, the workload (i.e., the request arrival rate) at the edge and the device $\lambda^{proc}_{dev}$ and $\lambda^{proc}_{edge}$ are identical. However, when an edge server is multiplexed across $m$ devices, its workload will be the sum of the individual device workloads under the offloading case, as shown in Figure 2. This higher workload will increase the queuing delays at the edge and may cause the end-to-end offloading latency to exceed the on-device processing latency.
Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process (Pishro-Nik, 2014), it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload. That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the $i-th$ device.
We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary. For this case, we model each system as an $M/G/1$ queue, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge. This extension captures the potentially large variability in service times across applications in multi-tenant environments. If $s_{i}$ denotes the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times. That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ and the the the effective service rate of the edge $\mu_{edge}=1/s_{edge}$. Then the aggregate utilization of the edge server $\rho_{edge}$ is $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$. This leads us to the following lemma for the multi-tenant case.
Lemma 3.0.
For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgeμedgeprocVar[sedge]2(μedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
Proof.
The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula (Gross et al., 2008; Harchol-Balter, 2013): {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (3.3) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into Equation 4 completes the proof. ∎
Practical takeaways. The expected wait time in an $M/G/1/FCFS$ system is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays. It is therefore important for application designers to consider service time variability when assigning applications to edge servers. Grouping applications with similar processing demands can help reduce overall variability, thereby minimizing queuing delays and improving edge offloading performance. Further, the above analysis assumes that isolation features such as GPU virtualization, which are available on server-grade GPUs, are not available on edge accelerators. If such features were available, each device could be assigned to an isolated virtual GPU on the edge. In this special case, the offloading situation is reduced to a situation similar to that considered in Sec 3.2.1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device. However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
4. Model Validation
In this section, we empirically evaluate how the key factors modeled in our analytic framework affect the performance of the two execution approaches and demonstrate that our models can accurately predict the response times across a wide range of configurations.
4.1. Experimental Setup
Workloads. We use DNN-based accelerator-driven workloads to validate our model’s ability to capture the impact of workload characteristics. We focus on cognitive assistant applications that use different DNN models for image classification and object detection. Specifically, we use three widely deployed DNN architectures for image classification and object detection, namely MobileNet, YOLO, and Inception. These models exhibit diverse computational demands and input dimensionalities as shown below:
Model Input Shape Params (M) FLOPs (G)
MobileNetV2 $[1,224,224,3]$ 3.5 0.6
InceptionV4 $[1,299,299,3]$ 42.7 6.3
YoloV8n $[1,640,640,3]$ 3.2 8.7
All inference tasks are executed using TensorFlow 2.13.0 across the evaluated hardware. We use frames extracted from a 720p dash cam video as input to the models, with each frame processed either locally on the device or offloaded to the edge. Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process. The workload generator runs on a separate machine to avoid performance interference with the device during evaluation.
Hardware. We use a diverse hardware platform to reflect real-world deployment scenarios. For mobile devices with accelerators, we use an NVIDIA Jetson TX2 and an NVIDIA Jetson Orin Nano, with TX2 being an earlier model and the Orin Nano a newer, more powerful model. For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU, a 40-core Intel Xeon E5-2660v3 CPU, and 256GB RAM. We also use a server with an NVIDIA RTX 4070 GPU, 8-Core AMD Ryzen 7 5700X CPU, and 32GB RAM. Both edge servers are connected to a 1 Gbps network. These hardware platforms span a wide range of performance levels, from the TX2’s 1.3 TFLOPS peak FP16 performance to the 2.6 TFLOPS of the Orin Nano, the 4.5 TFLOPS of the A2, and the 29.1 TFLOPS of the RTX 4070. For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable. To estimate $k$, we empirically measure how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior. All platforms run Ubuntu 18.04, and we utilize the Linux Traffic Control (TC) subsystem to emulate a range of network bandwidth in our experiments.
4.2. Parameter Estimation
Our model requires several parameters to predict the end-to-end response time for each execution strategy. Here, we describe how these parameters are obtained in our evaluation and outline other practical methods for estimating these parameters.
Estimating average service times. We estimate request service times by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool (NVIDIA, 2024), which reports per-process execution times on the accelerator. There is also extensive literature (Wang et al., 2024b; Yao et al., 2018; Kang et al., 2017; Wang et al., 2024b) on estimating the latency of neural network inference by analyzing factors such as network structure and parameter size, which achieves high accuracy due to the constant number of operations per inference across requests.
Estimating average network delays. Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf (The iPerf Project, 2025) to quickly adjust to changes. In addition to queuing and transmission delays captured by our analytic models, network delays also include propagation and processing delays. Since we assume that requests are offloaded to a nearby edge server, the propagation delay (caused by signals traveling at the speed of light) is minimal and thus excluded from our models. Processing delay can be estimated by measuring the round-trip time of small probe packets, and is excluded from our models due to its negligible impact on overall latency.
Estimating system utilization. The system utilization $\rho$ is calculated as the ratio of the arrival rate $\lambda$ to the service rate $\mu$ of the system. The arrival rate $\lambda$ can be estimated by applying a sliding window over incoming request timestamps. The service rate $\mu$ can be estimated from system logs based on the number of completed requests within a given time interval.
4.3. Impact of Workload Characteristics
Our first experiment investigates how workload characteristics influence offloading decisions. Specifically, we evaluate three workloads with distinct computational profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload due to its larger input dimensions). We compare on-device processing on TX2 and Orin Nano against offloading workloads to an NVIDIA A2 GPU, using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps. For each device, we test all available GPU frequency settings.
Figure 3. Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2, Right: Orin Nano vs A2.
Figure 2 compares the average latency of on-device processing across various GPU frequency settings with edge offloading under a 5 Mbps network. Figures 3(a) and 3(b) show that for MobileNetV2 (low compute, small payload), local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels. In contrast, the trend reverses for InceptionV4 (high compute, moderate payload), where offloading provides lower average latency across all device configurations, as shown in Figures 3(c) and 3(d). Figures 3(e) and 3(f) plot the average latencies for YOLOv8n (high compute, large payload). In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes, despite YOLOv8n’s high compute demand. Meanwhile, Orin Nano achieves lower average latency than offloading across all performance modes. We observe similar results when offloading to RTX4070, and omit the details here due to space constraints. In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%, proving the accuracy of our models.
Collectively, these empirical results align with the qualitative observations outlined in Remark 3.2 and Remark 3.3. While the performance gap between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands. For workloads with low computational demand (e.g., MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delay. However, for computationally intensive workloads (e.g., InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network costs of offloading. Moreover, workloads with large request payloads (e.g., YoloV8n) favor on-device processing because it does not involve any network components and can avoid involving a large network delay.
Key takeaway. Offloading provides lower latency when the edge’s absolute performance gain exceeds the associated network delay. Our models accurately predict the latency for both strategies with a mean absolute percentage error of 2.2%.
4.4. Impact of Network Conditions
Figure 4. Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
Figure 4 evaluates how network speed affects the effectiveness of each execution strategy. Figure LABEL:plot:mnet_varyNet_4070 compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds. For both devices, we observe a performance crossover point where offloading becomes faster than on-device processing. The location of the crossover depends on both network bandwidth and the compute performance gap between the device and the edge. For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial.
Figure LABEL:plot:mnet_varyNet_A2 further demonstrates the relationship between network speed and device-edge performance difference. We find that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance. For example, at 5 Mbps, the device outperforms edge offloading under its highest-performance mode. However, as network bandwidth improves to 10 Mbps, edge offloading becomes more favorable. Across all scenarios, our analytical models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.
Key takeaway. At high network bandwidths, edge offloading becomes favorable as its computational advantage outweighs the reduced network delays.
4.5. Impact of Request Rate
Figure 5. Latency comparison between on-device processing and offloading to A2 for MobileNetV2 with varying request rates under 10 Mbps and 20 Mbps networks.
Next, we demonstrate how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload. In this experiment, both the TX2 and Orin Nano are set to their maximum performance modes. We evaluate edge offloading under 10 Mbps and 20 Mbps network bandwidth.
Figure 5 plots the end-to-end latency of each strategy as a function of request rate (RPS). When comparing Orin Nano with edge offloading to A2, we observe two distinct behaviors depending on the network bandwidth. At 10 Mbps, on-device processing using Orin Nano outperforms edge offloading across all request rates. However, under a 20 Mbps network, there exists a performance crossover point: On-device processing on Orin Nano provides lower latency under low load (RPS ¡ 60), but edge offloading becomes the faster option once the RPS exceeds 60. On the other hand, TX2 consistently exhibits the highest latency across all request rates due to its limited processing capability. Overall, the predicted end-to-end latencies of our models match closely with the observed latency, proving the accuracy of our models.
This result shows that while both the network queueing delays (for offloading) and local processing queueing delay (for on-device processing) grow with $\lambda$, the rate at which delays increase differs depending on the hardware and network configurations. Under 10 Mbps, the network queueing delay increases at a faster rate than the device queueing delay as $\lambda$ increases, favoring on-device processing. Conversely, under 20 Mbps, device queuing delays dominate at higher $\lambda$, shifting the advantage to offloading. These competing scaling effects validate the observation we discussed in Lemma 3.1.
Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency under faster networks.
4.6. Impact of Multi-Tenancy
Figure 6. Latency comparison across execution strategies for InceptionV4 under a 100 Mbps network with varying number of co-located applications. Each application has RPS = 2.
We now experimentally validate our model’s ability to predict average response time in multi-tenant settings using Orin Nano as the device and A2 as the edge server. We vary the number of applications deployed on the edge, each serving InceptionV4 requests, with all models loaded in memory concurrently. Each application receives 2 requests per second from its corresponding client. We report the average offloading latency across all clients and compare it with on-device processing.
Figure 6 compares the end-to-end latency of offloading with on-device processing as the number of co-located applications increases. When only a single application is deployed on the edge, offloading yields lower latency than on-device processing. As the edge server hosts more applications, the latency of offloading increases due to resource contention across applications. We observe a performance crossover point at 18 co-located applications, beyond which on-device processing becomes more efficient. Note that the on-device processing time remains constant regardless of the number of concurrent applications as devices operate independently. In all cases, our models closely predict the latency for both execution strategies.
Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on the edge server.
5. Models in Action
In this section, we present a resource manager that leverages our analytic models to adaptively decide between offloading and on-device processing. We outline its decision algorithm and demonstrate its effectiveness through two case studies that showcase how the resource manager adapts to network variability and workload interference in multi-tenant edge environments.
5.1. Model-Driven Adaptive Resource Management
We implement a resource manager that utilizes our analytic models to adaptively switch between on-device and edge processing. The manager operates locally on the device and periodically collects runtime metrics such as network bandwidth, edge server load, and request arrival rate. For each epoch, it uses the models to estimate the average latency of both execution strategies and chooses the option with the lower predicted average latency to execute requests.
Algorithm 1 outlines the decision-making process for selecting the execution strategy in the resource manager. At Lines 1-2, it predicts on-device average execution latency based on the current arrival rate and device processing speed. At Lines 3-6, it computes the predicted average latency for offloading to each edge server, considering the server load, service rates, and the device’s network conditions. Finally, at Lines 7-11, it executes the request using the best execution strategy based on these predictions.
Input:
$D_{req},D_{res}$: Request/result payload size
$s_{dev}^{proc},s_{edge}^{proc}$: Service times of device and edge
$\lambda_{dev}$: Device task arrival rate
$B$: Network bandwidth of device
$\mathcal{E}$: Number of edge servers
$\{\mu_{edge,E}^{proc}\}_{E=1}^{\mathcal{E}}$: Edge server aggregated service rates
$\{\lambda_{edge,E}\}_{E=1}^{\mathcal{E}}$: Edge server aggregated arrival rates
Output:
Execution strategy (on-device or edge server $E^{*}$)
// Predict on-device processing latency
1 $\mu_{dev}^{proc}\leftarrow 1/s_{dev}^{proc}$
2 $T_{dev}\leftarrow M/D/1(\lambda^{proc}_{dev},\mu_{dev}^{proc})+s_{dev}^{proc}$
// Predict edge offloading latency
3 $T_{net}^{req}\leftarrow M/M/1(\lambda_{dev},B/D_{req})+D_{req}/B_{dev}$
4 $T_{net}^{res}\leftarrow M/M/1(\lambda_{edge,E},B/D_{res})+D_{res}/B_{edge}$
5 for $E\leftarrow 1$ to $\mathcal{E}$ do
6       $T_{edge,E}\leftarrow T_{net}^{req}+M/G/1(\lambda_{edge,E},\mu_{edge,E}^{proc})+s_{edge}^{proc}+T_{net}^{res}$
// Select optimal strategy
7 if $T_{dev}<\min(\{T_{edge,E}\}_{E=1}^{\mathcal{E}})$ then
8       return on_device_processing()
9else
10       $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
       return offload($E^{*}$) ;
        // Offload to server $E^{*}$
11
Algorithm 1 Model-Driven Adaptive Offloading Algorithm
5.2. Case 1: Fluctuating Network Conditions
Figure 7. Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
We first demonstrate how the resource manager adapts to changes in network conditions. We use the TX2 as the device, A2 as the edge server, and MobileNetV2 as the workload, with requests arriving at a fixed rate of 10 RPS. Initially, the device has a network bandwidth of 20 Mbps. To simulate network variation, we reduce the bandwidth to 10 Mbps at t = 20, further reduce it to 2 Mbps at t = 40, and then increase it back to 20 Mbps at t = 60.
Figure 7 shows the observed latency using the execution strategies selected by the resource manager, alongside the model-predicted latencies for both offloading and on-device processing. At t = 0, the predicted offloading latency is lower than on-device processing, therefore the manager chooses to offload requests. At t = 20, despite the reduced bandwidth from 20 to 10 Mbps, the predicted latency for offloading remains lower, and the strategy remains unchanged. However, at t = 40, the drop to 2 Mbps causes the predicted offloading latency (45 ms) to exceed the predicted on-device processing time (21 ms), and the resource manager to switch to on-device processing. Finally, at t = 60, as the network returns to 20 Mbps, the resource manager switches back to offloading. Throughout this sequence, the models closely predict the observed values, validating their accuracy.
Key takeaway. Our resource manager dynamically adapts to changing network conditions by switching between execution strategies.
5.3. Case 2: Multi-tenant Edge Servers
Figure 8. Observed latency under varying edge server workloads. At t = 80, the request rate of edge server 1 is increased from 10 RPS to 50 RPS. At t = 160, the request rate of edge server 2 is also increased to 50 RPS.
Next, we demonstrate how our resource manager adapts to workload changes across multiple multi-tenant edge servers. In this experiment, two edge servers, $E_{1}$ and $E_{2}$, each equipped with an A2 GPU, are hosting separate instances of a YOLOv8n application. Initially, the request rates are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while a TX2 running the resource manager has a request rate of $\lambda_{dev}=10$. At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50. Similarly, at t = 160, we increase $\lambda_{E_{2}}$ to 50.
Figure 8 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing. Initially, from t = 0 to t = 80, the predicted latency for offloading to $E_{1}$ is the lowest, therefore the resource manager offloads requests to $E_{1}$. However, at t = 80, as the workload on $E_{1}$ increases, its predicted latency rises. In response, the resource manager adapts by offloading to the less loaded server $E_{2}$. At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and it shifts to on-device processing. This sequence demonstrates the resource manager’s ability to dynamically adapt to fluctuations in edge server workload.
Key takeaway. Our resource manager dynamically adapts to workload changes in multi-tenant edge servers to maintain low latency.
6. Related Work
Edge offloading. Edge offloading has become a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure (Davis et al., 2004; Jang et al., 2018; Lai et al., 2017; Li et al., 2024; Ogden et al., 2021; Ren et al., 2019; Mei et al., 2012; Liu et al., 2018; Chen et al., 2017; Wang et al., 2018; Zhang et al., 2024a; Tkachenko et al., 2023). For instance, Furion (Lai et al., 2017) offloads computation-intensive components of mobile AR pipelines to the edge. Jang et al.(Jang et al., 2018) propose an edge camera virtualization architecture to offload video analytics workloads. Li et al.(Li et al., 2024) introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints. In contrast, our work uses a model-driven approach to reexamine whether offloading is advantageous in the era of accelerators. We extend our analysis to capture the combined effects of hardware characteristics, workload and network dynamics, and multi-tenancy in edge servers.
Model-aware placement and resource allocation. Queueing models have been extensively used to predict the performance of traditional CPU-bound workloads (Jiang et al., 2021; Wang et al., 2024a; Tantawi and Towsley, 1985; Ambati et al., 2020; Gandhi and Suresh, 2019; Shen et al., 2011; Urgaonkar et al., 2005). Recently, there has been work focusing on modeling the performance of accelerators. For example, Reference (Merck et al., 2019) models the latency of GPU batching in DNN applications. Liang et al. (Liang et al., 2023) model the performance of executing inference requests on various accelerators (e.g., TPU, GPU) and propose an online knapsack algorithm that uses the models for DNN model placement. In contrast, our models incorporate not only device hardware characteristics but also network delays and device-to-edge performance differences, enabling direct comparisons between on-device and edge offloading strategies.
Collaborative inference. There is a growing body of work on collaborative inference, where the computation is split across multiple nodes in the network to reduce end-to-end latency (Almeida et al., 2022; Banitalebi-Dehkordi et al., 2021; Gao et al., 2022; Hu et al., 2019; Huang et al., 2020; Kang et al., 2017; Laskaridis et al., 2020; Li et al., 2018; Merck et al., 2019; Ng et al., 2024; Ogden and Guo, 2023; Tuli et al., 2023; Yang et al., 2022; Zhang et al., 2020, 2024b). Neurosurgeon (Kang et al., 2017) proposes regression models for layer-wise execution times prediction for DNN partitioning across network components. SPINN (Laskaridis et al., 2020) integrates early-exit networks with model partitioning across device and cloud, enabling graceful degradation under network instability. In this work, we focus on offloading entire inference requests and do not consider partitioned execution.
7. Conclusions
In this paper, we presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators. We validated the accuracy of our model across a range of hardware, workloads, and network configurations. We developed a resource manager that leverages these predictions to dynamically adapt to changing network conditions and multi-tenant edge server workloads. As future work, we plan to extend our models to incorporate collaborative processing, i.e., performing a portion of request computation locally before offloading.

### v2

@@
 To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing
-Abstract.
+Abstract
 Computational offloading is a promising approach for overcoming resource constraints on client devices by moving some or all of an application’s computations to remote servers.
@@
 We use our models to develop an adaptive resource manager for intelligent offloading and show its efficacy in the presence of variable network conditions and dynamic multi-tenant edge settings.
-submissionid: 696copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXX
-1. Introduction
+publicationid: pubid:
+I Introduction
 Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment.
 In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements.
-Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference (Mosqueira-Rey et al., 2022).
-Edge computing, which is a complementary approach to cloud computing, is a promising method for meeting the needs of such latency-sensitive applications.
+Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference [1].
+Edge computing, which is a complementary approach to cloud computing, is a promising method to meet the needs of such latency-sensitive applications.
 Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing.
-However, recent research (Ali-Eldin et al., 2021; Wang et al., 2024a) has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud.
+However, recent research [2, 3] has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud.
 In particular, since edge clusters are resource-constrained relative to the cloud, research has shown resource bottlenecks can arise in the presence of time-varying workloads and increase the processing latency at the edge, negating its network latency advantage.
-Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing (Wang et al., 2024a).
+Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing [3].
 A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge.
 [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper.
-Computational offloading (Satyanarayanan et al., 2009; Davis et al., 2004) involves using nearby edge resources to take on some, or all, of an application’s processing demands.
+Computational offloading [4, 5] involves using nearby edge resources to take on some, or all, of an application’s processing demands.
 Such offloading is performed to overcome resource constraints on a local device by leveraging more abundant processing capacity at edge servers.
-Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers (Ren et al., 2019).
-In this case, local devices pay a so-called mobility penalty (Satyanarayanan et al., 2019), which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead.
-Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks (Jang et al., 2018).
+Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers [6].
+In this case, local devices pay a so-called mobility penalty [7], which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead.
+Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks [8].
 In these offloading scenarios, conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
 With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought.
-For example, the rise of the so-called AI PC with onboard AI accelerators (Intel Corporation, 2025) has enabled local AI processing for many applications without a need for edge offloading.
-Similarly, mobile phones are equipped with neural accelerators (Simonite, 2017) that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally (Netatmo, 2025).
+For example, the rise of the so-called AI PC with onboard AI accelerators [9] has enabled local AI processing for many applications without a need for edge offloading.
+Similarly, mobile phones are equipped with neural accelerators [10] that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally [11].
 At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before.
@@
 Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators.
-Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and the choice of where to perform the processing will depend on hardware, network, and workload characteristics.
+Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and that the choice of where to perform the processing depends on hardware, network, and workload characteristics.
 In developing our analytic models and validating this hypothesis, this paper makes the following contributions.
   • We use queuing theory-based models to model the behavior of device and edge accelerators when processing requests for local on-device processing and edge offloading scenarios.
+Our modeling follows a two-level approach: it takes estimated service times obtained through profiling or other prediction techniques as input and embeds them into queuing models to obtain end-to-end response time predictions.
 Using these models, we develop closed-form analytic bounds to determine when one approach outperforms the other under different hardware, network, and workload scenarios.
-We also develop analytic models for comparing on-device processing to multi-tenant edge servers that service multiple clients.
-  • We experimentally validate our analytic models and bounds using several different device and edge accelerators and network configurations and show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios.
-Our evaluation shows that our model achieves a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
-  • We develop a dynamic resource manager that incorporates our models and demonstrate the ability of this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios.
+We also develop analytic models for comparing on-device processing to multi-tenant edge servers that serve multiple clients.
+Although our primary focus is on offloading, we also show how the models can be extended to account for device-edge collaborative (i.e., split) processing.
+  • We experimentally validate our analytic models and bounds using a variety of device and edge accelerators, network configurations, and workloads including deep neural networks (DNNs) and recurrent neural networks (RNNs).
+Our results show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
+  • We develop a dynamic resource manager that incorporates our models and demonstrate how they can be leveraged by this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios.
 Our results show that it effectively adapts to changing network conditions and fluctuating request rates at edge servers.
-2. Background
+II Background
 This section provides background on computational offloading, hardware accelerators, and on-device processing.
-Table 1. Characteristics of common programmable accelerators.
+TABLE I: Characteristics of common programmable accelerators.
 Accelerator Peak Power (W) Memory Cost Target Workload
@@
 NVIDIA A2 GPU 60 16 GB of GDDR6 \$1399 AI and GPU workload
-2.1. Cloud and Edge Offloading
+II-A Cloud and Edge Offloading
 Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server.
-In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks (Satyanarayanan et al., 2009; Davis et al., 2004).
-While such edge offloading involves a network hop to the edge server (also known as the mobility penalty (Satyanarayanan et al., 2019)), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices.
-In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload (Huang et al., 2020).
+In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks [4, 5].
+While such edge offloading involves a network hop to the edge server (also known as the mobility penalty [7]), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices.
+In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload [12].
 For example, large language model (LLM) inference uses GPU clusters in the cloud to answer user queries, while IoT devices use cloud storage and processing to perform analytic and data processing tasks.
 In general, computational offloading offers benefits when local processing is constrained and the cost of incurring the additional network latency outweighs the benefits of accessing faster processing in the edge or the cloud.
-2.2. Hardware Accelerators
+II-B Hardware Accelerators
 In recent years, a wide range of programmable accelerators has been developed to accelerate a range of compute tasks, particularly machine learning compute.
@@
 Table 1 depicts the characteristics of several commonly deployed device and edge accelerators, ranging from low-end to high-end options.
-Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications and CNN inference (Intel, 2017).
-Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge (Google, 2019).
-NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision (NVIDIA, 2017, 2023, 2021).
+Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications [13].
+Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge [14].
+NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision [15, 16, 17].
 For these specific workloads, programmable accelerators can provide substantial performance improvements, delivering server-like capabilities for some compute-intensive tasks even on low-end devices.
 With their small form factor and energy-efficient design, integrating accelerators into mobile devices has become increasingly common.
-For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device (Simonite, 2017).
+For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device [10].
 With the presence of accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without the need to resort to edge offloading.
-At the same time, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks.
+Meanwhile, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks.
 This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
-2.3. On-device Processing versus Edge Offloading
+II-C On-device Processing versus Edge Offloading
 As noted above, the availability of accelerators on client devices has enabled on-device processing of many tasks without the need to offload those tasks to the edge or to the cloud.
@@
 On the one hand, mobile devices are subject to network variability due to fluctuating wireless signal strength, interference, and congestion in shared communication channels.
-Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting the overall offloading performance.
+Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting its overall performance.
 Additionally, mobile devices running applications with prolonged operating durations, such as live object detection on captured video, may experience battery constraints.
@@
 Moreover, edge servers are commonly multiplexed across different applications to maximize edge resources utilization.
-At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs (NVIDIA, 2025), leading to performance interference between applications.
+At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs [18], leading to performance interference between applications.
 As a result, execution times can become variable, making it challenging to determine whether to offload under dynamic workloads.
@@
 To address this question, the next section develops analytic models to capture the end-to-end execution time for each on-device and edge processing, capturing the key factors that impact application performance.
-3. Model-driven Performance Comparison
-In this section, we develop analytical queuing models for edge offloading and on-device processing and then develop closed-form expressions for their expected latencies to compare the two approaches and determine when one strategy outperforms the other.
-3.1. Analytic Queuing Models
-(a) Edge offloading queueing model
-(b) On-device processing queueing model
-Figure 1. Modeling request execution using (a) edge offloading and (b) on-device processing.
+III Model-driven Performance Comparison
+In this section, we develop analytical queuing models for edge offloading and on-device processing, and derive closed-form expressions for their expected latencies to compare the two approaches and identify when one outperforms the other.
+III-A Analytic Queuing Models
+(a) Edge offloading queuing model
+(b) On-device processing queuing model
+Figure 1: Modeling request execution using (a) edge offloading and (b) on-device processing.
 Our analytical models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems.
-Queuing theory has previously been used to model the performance of edge and cloud applications (Ali-Eldin et al., 2021; Harchol-Balter, 2013; Urgaonkar et al., 2005), and will serve as the foundation for analyzing edge offloading and on-device processing.
+Queuing theory has previously been used to model the performance of edge and cloud applications [2, 19, 20], and will serve as the foundation for analyzing edge offloading and on-device processing.
 As discussed below, we model the end-to-end latency of a request—the duration of time from when the task enters the first queue until it exits the system—for both edge offloading and on-device processing using queuing theory.
@@
 To model the end-to-end request execution time using edge offloading, we need to model the times spent by the request both at the device and at the edge server.
-We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1(a).
+We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1a.
 We assume that the request arrives at the device and is then forwarded (“offloaded”) to the edge server over the network.
@@
 On-device processing.
-Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1(b).
+Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1b.
 Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores of the device where they are processed with a degree of parallelism $k_{dev}$.
@@
 Also, let $n_{req}$ denote the network delay of sending a request from the device to the edge, and $n_{res}$ be the network delay of sending a result from the edge to the device.
-With the above notations, we now model the end-to-end latency of requests executed using each strategy. First, we have
-(1) $T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
+With the above notations, we now model the end-to-end latency of requests executed using each strategy:
+$T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (1)
 and
-(2) $T_{dev}=w_{dev}^{proc}+s_{dev}$
-where $T_{edge}$ and $T_{dev}$ denote the end-to-end latency of requests executed using edge offloading and on-device processing, respectively.
-3.2. Deriving and Analyzing End-to-End Latency
+$T_{dev}=w_{dev}^{proc}+s_{dev}$ (2)
+where $T_{edge}$ and $T_{dev}$ denote the end-to-end request latency under edge offloading and on-device processing, respectively.
+III-B Deriving Service Times on Accelerators
+Our analytic models require the service time of a workload as input to compute end-to-end response times.
+The service time reflects the computational demand on the target accelerator and indicates how quickly a request is processed on a given platform.
+It can be obtained through empirical profiling or a deep learning model that predicts workload execution times on accelerators.
+For empirical profiling, many applications already support function-level performance logging or provide profiling hooks [21], which can be used to derive the service time of a request by analyzing the corresponding function’s execution metrics.
+Alternatively, the workload can be profiled on the target hardware using benchmarking tools [22, 23].
+In case of model-based prediction, a neural network can be trained to predict the service time of a model on a given target hardware.
+For example, prior work [24, 25, 26] estimates DNN inference latency by analyzing factors such as network structure and parameter size.
+For recurrent architectures such as Gated Recurrent Unit (GRU) networks and Long Short-Term Memory (LSTM) networks, existing latency estimation techniques [27, 28] incorporate factors such as temporal dependencies, hidden state dimensionality, and sequence unrolling behavior to accurately predict inference latency.
+While any of these prior techniques can be used with our analytic models, our experiments use a simple neural network model to predict the service times.
+III-C Deriving and Analyzing End-to-End Latency
 In this section we compute the expected latency for edge offloading and on-device processing.
-Our analysis focuses on applications that require processing results on the device, such as AR/VR applications.
-Note that the proposed analytic models can also capture the performance of applications requiring results on the edge, and our findings can be extended to these cases.
+Our analysis focuses on applications that require processing results on the device such as AR/VR applications.
+Note that our models can also be extended to capture the performance of applications that require results on the edge.
 In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
-3.2.1. Comparing Edge Offload and On-Device Execution
+III-C1 Comparing Edge Offload and On-Device Execution
 We begin our analysis by examining the conditions under which edge offloading results in higher latency compared to on-device processing.
-For workloads that benefit from hardware acceleration, such as those utilizing GPUs or TPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs.
+For workloads that benefit from hardware acceleration such as those utilizing GPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs.
 In the following discussion, we refer to such workloads as accelerator-driven workloads.
-Table 2 summarizes the notation used in our models.
-Table 2. Notations used in analytic models.
+While our analysis primarily focuses on DNN inference, the results generalize to other AI workloads as will be discussed in Section III-E.
+Table II summarizes the notation used in our models.
+TABLE II: Notations used in analytic models.
 Notation Description
@@
 $\mu^{proc}_{dev},\mu^{proc}_{edge}$ Service rate of device/edge
-Lemma 3.0.
-For accelerator-driven workloads, edge offloading incurs a larger average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
+Lemma III.1.
+For accelerator-driven workloads, edge offloading incurs a higher average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
 Proof.
 Edge offloading will offer a worse latency than on-device processing when $T_{device}<T_{edge}$. That is
-(3) $w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
+$w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (3)
 Let $D_{req}$ and $D_{res}$ denote the task and result payload size, respectively. Further, let $B$ denote the bandwidth of the device-to-edge network path. The above inequality reduces to
-(4) $s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$
-As shown in (Liang et al., 2023), workloads on programmable accelerators typically exhibit deterministic service times (denoted as $D$), as the number of operations per request remains constant.
-Since typical accelerators have a small degree of parallelism $k$ (Amert et al., 2017), the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process.
-While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ (Shah et al., 2018, 2020).
+$s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$ (4)
+As shown in [29], DNN inference on programmable accelerators often exhibits deterministic service times (denoted as $D$), as the number of operations per request remains constant.
+Since typical accelerators have a small degree of parallelism $k$ [30], the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process.
+While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ [31, 32].
 This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$.
-Its expected wait time can be derived from the well-known Pollaczek-Khinchin (P-K) formula (Gross et al., 2008; Harchol-Balter, 2013) and is given by
-(5) $E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$
+Its expected wait time can be derived from the Pollaczek-Khinchin (P-K) formula [33, 19] and is given by
+$E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$ (5)
 On the other hand, network interfaces with a single controller can be modeled as $M/M/1/FCFS$ systems with expected wait times given by
-(6) $E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$
+$E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$ (6)
 Let $k_{dev}$ and $k_{edge}$ denote the degree of parallelism seen at the accelerator on the device and the edge respectively.
@@
 Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$.
-Substituting the above into (3.2.1), we obtain: {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof.
+Substituting the above into (III-C1), we obtain {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof.
 ∎
 The above result implies the following two remarks.
-Remark 3.2.
+Remark III.1.
 On-device processing is likely to outperform edge offloading for workloads with lower computational demand.
@@
 For requests with lower computational demands, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally due to the disparity in processing capacities between the device and the edge as the workload computational demand decreases.
-Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (3.1) will also become smaller since $\mu$ increases for lighter requests.
+Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (III.1) will also become smaller since $\mu$ increases for lighter requests.
 The terms representing the network overheads remain unchanged since they depend on request size rather than the processing demand of the request.
-As a result, the inequality in (3.1) is more likely to hold true for requests with lighter processing demand.
+As a result, the inequality in (III.1) is more likely to hold true for requests with lighter processing demand.
 ∎
-Remark 3.3.
-On-device processing is more likely to outperform edge offloading for slower networks or when workloads have larger request or result payloads.
+Remark III.2.
+On-device processing is likely to outperform edge offloading on slower networks or when workloads have larger request or result payloads.
 Explanation.
-The term $\frac{D_{req}+D_{res}}{B}$ in (3.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality.
+The term $\frac{D_{req}+D_{res}}{B}$ in (III.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality.
 In addition, longer transmission times reduce the service rates of the NIC at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the RHS.
@@
 Practical takeaways.
-Lemma 3.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective.
+Lemma III.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective.
 The bound and the resulting inequality depends on several factors: (i) relative processing speeds of the device and edge, denoted by $s_{dev}$, $s_{edge}$ (and also the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) network speed, (iii) overall workload denoted by request rate $\lambda$, (iv) network transmission overheads.
@@
 Thus, the optimal strategy depends on the workload’s payload size relative to both network bandwidth and the edge-device performance gap.
-Remark 3.2 shows how the computational demand of the workload influences the execution strategy.
+Remark III.1 shows how the computational demand of the workload influences the execution strategy.
 For workloads with higher computational demand, edge offloading is more likely to outperform on-device processing, as the edge’s processing speed advantage becomes larger.
 Conversely, for workloads with lower computational demand, on-device processing is more efficient as the performance advantage of edge processing is likely outweighed by the associated network costs.
-Remark 3.3 demonstrates how network delays affect the efficiency of edge offloading.
+Remark III.2 demonstrates how network delays affect the efficiency of edge offloading.
 With smaller network bandwidth or larger request/result payloads, both the network queuing delay and transmission time increase.
-As a result, edge offloading becomes less efficient, while on-device processing is unaffected since processing is local and does not rely on network resources.
-3.3. Impact of Multi-Tenancy
-Figure 2. $m$ devices offloading requests to an edge server.
-While the analysis in the Sec 3.2.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work.
+As a result, edge offloading becomes less efficient, while on-device processing is unaffected since it does not rely on network resources.
+Extending to collaborative processing.
+Collaborative (i.e., split) processing involves partially processing requests locally on the client’s device before offloading them to a remote edge server to minimize the amount of data sent over the network [34, 35, 36, 12, 26, 37, 38].
+To compute the end-to-end response time of such a strategy, we combine our analytic models for on-device processing and edge offloading into a tandem queuing network that accounts for both the partial local computation and the subsequent transmission and processing delays at the edge server.
+Specifically, a request first enters the components modeled in Figure 1b and is partially processed with a service time $s^{\prime}_{dev}$ representing the partial local processing time.
+Then, the intermediate output of size $D_{inter}$ passes through the components modeled in Figure 1a, with a network transmission time of $D_{inter}/B$ and a service time $s^{\prime}_{edge}$ representing the remaining computation time on the edge.
+This combined model can then be used to estimate performance and inform decision-making when selecting between local, edge, or collaborative processing strategies.
+III-D Impact of Multi-Tenancy
+Figure 2: $m$ devices offloading requests to an edge server.
+While the analysis in Lemma III.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work.
 Hence, we extend our model-driven comparison to multi-tenancy scenarios.
@@
 This higher workload will increase the queuing delays at the edge and may cause the end-to-end offloading latency to exceed the on-device processing latency.
-Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process (Pishro-Nik, 2014), it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload.
+Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process [39], it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload.
 That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the $i-th$ device.
 We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary.
-For this case, we model each system as an $M/G/1$ queue, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge.
+For this case, we model each system as an $M/G/1$ system, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge.
 This extension captures the potentially large variability in service times across applications in multi-tenant environments.
@@
 This leads us to the following lemma for the multi-tenant case.
-Lemma 3.0.
-For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgeμedgeprocVar[sedge]2(μedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
+Lemma III.2.
+For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgekedgeμedgeprocVar[sedge]2(kedgeμedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
 Proof.
-The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula (Gross et al., 2008; Harchol-Balter, 2013): {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (3.3) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into Equation 4 completes the proof.
+The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula [33, 19] {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (III-D) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into (4) completes the proof.
 ∎
 Practical takeaways.
-The expected wait time in an $M/G/1/FCFS$ system is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays.
+The expected wait time in $M/G/1/FCFS$ systems is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays.
 It is therefore important for application designers to consider service time variability when assigning applications to edge servers.
@@
 If such features were available, each device could be assigned to an isolated virtual GPU on the edge.
-In this special case, the offloading situation is reduced to a situation similar to that considered in Sec 3.2.1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device.
+In this special case, the offloading situation is reduced to a situation similar to that considered in Sec III-C1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device.
 However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
-4. Model Validation
+III-E Discussion
+Generalizing to other accelerator-driven workloads.
+While the above analysis focuses on DNN workloads with deterministic service times modeled using an $M/D/1$ system, our modeling framework can be extended to other workloads with different service time characteristics by incorporating appropriate queuing results.
+For example, workloads with recurrent architectures such as GRUs and LSTMs often process variable-length inputs with service times that vary with the input sequence length.
+Similarly, early-exit DNNs allow requests to exit at different layers based on confidence thresholds, which results in variable service times.
+To capture such variability, our framework can adopt an $M/M/1$ queuing system where service times follow an exponential distribution with the expected wait time shown in (6).
+By selecting the appropriate queuing model, our framework can adapt to diverse workloads and provide accurate response time estimates as will be shown in Section IV-D.
+Currently, our framework does not account for workloads that violate the FCFS assumption, such as Large Language Model (LLM) inference with continuous batching where requests are dynamically batched, and extending the models to capture such behavior is a subject of future work.
+Model limitations.
+Our analysis assumes request arrivals follow a Poisson process and are independent, which may not hold under bursty arrivals.
+To address this, our models can adopt a $G/G/1$ system, which accommodates arbitrary arrival and service time distributions.
+Although no closed-form solution exists for the expected wait time in this setting, Marshall et al.
+[40] provide an upper bound {dmath} E[w_G/G/1] ≤λ(σa2+ σs2)2(1-ρ) where $\sigma_{a}$ and $\sigma_{s}$ are the variances for the interarrival and service distributions respectively.
+Additionally, the queuing results used in our analysis assume that each hardware queue has infinite capacity.
+For scenarios with bounded queue sizes, we can incorporate results for finite-buffer queuing systems [19] to model the impact of such capacity constraints.
+IV Model Validation
 In this section, we empirically evaluate how the key factors modeled in our analytic framework affect the performance of the two execution approaches and demonstrate that our models can accurately predict the response times across a wide range of configurations.
-4.1. Experimental Setup
+IV-A Experimental Setup
 Workloads.
-We use DNN-based accelerator-driven workloads to validate our model’s ability to capture the impact of workload characteristics.
-We focus on cognitive assistant applications that use different DNN models for image classification and object detection.
-Specifically, we use three widely deployed DNN architectures for image classification and object detection, namely MobileNet, YOLO, and Inception.
-These models exhibit diverse computational demands and input dimensionalities as shown below:
+We validate our model’s ability to capture the impact of workload characteristics using a diverse set of workloads, including DNN-based models as well as recurrent models.
+For DNN workloads, we focus on cognitive assistant applications involving image classification and object detection.
+Specifically, we use three widely deployed DNN architectures, namely MobileNet, Inception, and YOLO, each with distinct computational demands and input sizes as shown below:
 Model Input Shape Params (M) FLOPs (G)
-MobileNetV2 $[1,224,224,3]$ 3.5 0.6
-InceptionV4 $[1,299,299,3]$ 42.7 6.3
-YoloV8n $[1,640,640,3]$ 3.2 8.7
-All inference tasks are executed using TensorFlow 2.13.0 across the evaluated hardware.
+MobileNetV2 $[224,224,3]$ 3.5 0.6
+InceptionV4 $[299,299,3]$ 42.7 6.3
+YOLOv8n $[640,640,3]$ 3.2 8.7
+All inference requests are executed using TensorFlow 2.13.0 across the evaluated hardware.
 We use frames extracted from a 720p dash cam video as input to the models, with each frame processed either locally on the device or offloaded to the edge.
-Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process.
-The workload generator runs on a separate machine to avoid performance interference with the device during evaluation.
+Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process, which runs on a separate machine to avoid performance interference with the device during evaluation.
+For recurrent models, we focus on personal voice assistant applications involving sentiment classification of conversational utterances.
+We validate our models using two types of recurrent neural networks: a GRU model with lower computational demand and a more compute-intensive LSTM model, with their architectural specifications summarized below:
+Model Embedding Dim # Layers $\times$ Units Activation
+GRU 512 2 $\times$ 512 Sigmoid
+LSTM 1024 3 $\times$ 1024 Sigmoid
+Both models are implemented using TensorFlow’s Keras API.
+To simulate real-world user inputs, we use the Amazon Alexa Topical Chat dataset [41], a knowledge-grounded human-to-human conversation dataset consisting of spoken-style dialogues with an average utterance length of 19 tokens.
 Hardware.
-We use a diverse hardware platform to reflect real-world deployment scenarios.
-For mobile devices with accelerators, we use an NVIDIA Jetson TX2 and an NVIDIA Jetson Orin Nano, with TX2 being an earlier model and the Orin Nano a newer, more powerful model.
+We use a diverse set of hardware platforms to reflect real-world deployment scenarios.
+For client devices with accelerators, we use the NVIDIA Jetson TX2 and Jetson Orin Nano, two commonly deployed platforms in IoT applications, with the TX2 being an earlier model and the Orin Nano offering better performance.
 For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU, a 40-core Intel Xeon E5-2660v3 CPU, and 256GB RAM.
@@
 Both edge servers are connected to a 1 Gbps network.
-These hardware platforms span a wide range of performance levels, from the TX2’s 1.3 TFLOPS peak FP16 performance to the 2.6 TFLOPS of the Orin Nano, the 4.5 TFLOPS of the A2, and the 29.1 TFLOPS of the RTX 4070.
+These platforms span a wide performance range, from 1.3 TFLOPS (TX2) to 2.6 TFLOPS (Orin Nano), 4.5 TFLOPS (A2), and up to 29.1 TFLOPS (RTX 4070) in peak FP16 throughput.
 For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable.
 To estimate $k$, we empirically measure how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior.
-All platforms run Ubuntu 18.04, and we utilize the Linux Traffic Control (TC) subsystem to emulate a range of network bandwidth in our experiments.
-4.2. Parameter Estimation
-Our model requires several parameters to predict the end-to-end response time for each execution strategy.
-Here, we describe how these parameters are obtained in our evaluation and outline other practical methods for estimating these parameters.
+All platforms run Ubuntu 18.04, and we use the Linux Traffic Control (TC) subsystem to emulate different network bandwidths.
+IV-B Parameter Estimation
+Our model requires several parameters to compute the end-to-end response times.
+We now describe how these parameters are obtained in our evaluation and outline other practical methods for estimating them.
 Estimating average service times.
-We estimate request service times by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool (NVIDIA, 2024), which reports per-process execution times on the accelerator.
-There is also extensive literature (Wang et al., 2024b; Yao et al., 2018; Kang et al., 2017; Wang et al., 2024b) on estimating the latency of neural network inference by analyzing factors such as network structure and parameter size, which achieves high accuracy due to the constant number of operations per inference across requests.
+Our system estimates workload service times through empirical profiling and a neural network–based prediction approach.
+We estimate service times for DNN workloads by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool [23], which reports per-process execution times on the accelerator.
+For GRU and LSTM workloads, we profile request latencies on a representative input set distinct from the test set and use the average latency from this profiling as the service time of the model.
+This profiling data is used for our edge-offloading and on-device processing experiments.
+For collaborative processing, a neural network is trained using profiling data to predict service time based on prior work discussed in Section III-B.
+Such a neural network approach avoids the extensive profiling overhead of all possible split configurations by predicting partial execution service times.
 Estimating average network delays.
-Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf (The iPerf Project, 2025) to quickly adjust to changes.
-In addition to queuing and transmission delays captured by our analytic models, network delays also include propagation and processing delays.
+Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf [42].
+In addition to queuing and transmission delays captured by our models, network delays also include propagation and processing delays.
 Since we assume that requests are offloaded to a nearby edge server, the propagation delay (caused by signals traveling at the speed of light) is minimal and thus excluded from our models.
@@
 The service rate $\mu$ can be estimated from system logs based on the number of completed requests within a given time interval.
-4.3. Impact of Workload Characteristics
-Our first experiment investigates how workload characteristics influence offloading decisions.
-Specifically, we evaluate three workloads with distinct computational profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload due to its larger input dimensions).
-We compare on-device processing on TX2 and Orin Nano against offloading workloads to an NVIDIA A2 GPU, using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps.
-For each device, we test all available GPU frequency settings.
-Figure 3. Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2, Right: Orin Nano vs A2.
-Figure 2 compares the average latency of on-device processing across various GPU frequency settings with edge offloading under a 5 Mbps network.
-Figures 3(a) and 3(b) show that for MobileNetV2 (low compute, small payload), local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels.
-In contrast, the trend reverses for InceptionV4 (high compute, moderate payload), where offloading provides lower average latency across all device configurations, as shown in Figures 3(c) and 3(d).
-Figures 3(e) and 3(f) plot the average latencies for YOLOv8n (high compute, large payload).
-In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes, despite YOLOv8n’s high compute demand.
-Meanwhile, Orin Nano achieves lower average latency than offloading across all performance modes.
-We observe similar results when offloading to RTX4070, and omit the details here due to space constraints.
+IV-C Impact of Workload Characteristics
+Our first experiment investigates how workload characteristics influence offloading decisions using three DNN workloads with distinct profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload).
+We compare on-device processing on TX2 and Orin Nano against offloading to an A2 GPU using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps.
+For each device, we test all available GPU frequency settings to emulate different device conditions.
+Figure 3: Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2; Right: Orin Nano vs A2.
+Figure 3 compares the average latency of on-device processing and edge offloading across workloads, devices, and GPU frequency levels.
+Figures 3a and 3b show that for MobileNetV2, local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels.
+The trend reverses for InceptionV4 as shown in Figures 3c and 3d, where offloading provides lower average latency across all device configurations.
+Lastly, Figures 3e and 3f shows the results for YOLOv8n.
+In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes despite YOLOv8n’s high compute demand, while Orin Nano achieves lower average latency than offloading across all performance modes.
+We observe similar results when offloading to RTX4070 and omit the details here due to space constraints.
 In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%, proving the accuracy of our models.
-Collectively, these empirical results align with the qualitative observations outlined in Remark 3.2 and Remark 3.3.
-While the performance gap between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands.
-For workloads with low computational demand (e.g., MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delay.
-However, for computationally intensive workloads (e.g., InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network costs of offloading.
-Moreover, workloads with large request payloads (e.g., YoloV8n) favor on-device processing because it does not involve any network components and can avoid involving a large network delay.
+Collectively, these empirical results align with the qualitative observations stated in Remark III.1 and Remark III.2.
+While the performance ratio between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands.
+For workloads with low computational demand (MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delays of offloading.
+However, for compute-intensive workloads (InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network delays.
+Moreover, workloads with large message payloads (YOLOv8n) favor on-device processing because it does not involve any network components and thus avoids any network delay.
 Key takeaway.
@@
 Our models accurately predict the latency for both strategies with a mean absolute percentage error of 2.2%.
-4.4. Impact of Network Conditions
-Figure 4. Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
-Figure 4 evaluates how network speed affects the effectiveness of each execution strategy.
-Figure LABEL:plot:mnet_varyNet_4070 compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds.
+IV-D Impact of Complex Deep Learning Models
+Figure 4: Latency comparison across execution strategies for (a) GRU model and (b) LSTM model at 3 RPS under 10 Mbps.
+Figure 5: Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
+Figure 6: Latency comparisons across (a) different collaborative processing configurations, (b) request rates under 10 Mbps and 20 Mbps networks, and (c) numbers of co-located applications.
+In addition to the DNN workloads, our model also accurately captures the latency behavior of more complex deep learning workloads.
+Figure 4 compares the average latency of on-device processing on the Orin Nano against offloading to A2 for two recurrent models: GRU and LSTM, under a moderate load of 3 RPS.
+Since service times vary with input length in these models, we incorporate the $M/M/1$ queuing model into our analysis to account for this variability in latency prediction.
+Figure 4a shows that even for the lightweight GRU model, offloading offers lower latency due to the small input utterance length (on average less than 100 bytes), which incurs minimal network delay.
+Figure 4b shows that as the model complexity increases, the absolute performance gain of offloading increases, leading to a greater latency reduction.
+In both cases, our model predictions closely match the observed latencies.
+Key takeaway. Our models generalize to complex deep learning workloads beyond DNN inference by integrating appropriate queuing models.
+IV-E Impact of Network Conditions
+Figure 5 shows how network speed affects the effectiveness of each execution strategy.
+Figure 5a compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds.
 For both devices, we observe a performance crossover point where offloading becomes faster than on-device processing.
-The location of the crossover depends on both network bandwidth and the compute performance gap between the device and the edge.
+The location of the crossover depends on both network bandwidth and the performance gap between the device and the edge.
 For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial.
-Figure LABEL:plot:mnet_varyNet_A2 further demonstrates the relationship between network speed and device-edge performance difference.
-We find that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance.
-For example, at 5 Mbps, the device outperforms edge offloading under its highest-performance mode.
-However, as network bandwidth improves to 10 Mbps, edge offloading becomes more favorable.
-Across all scenarios, our analytical models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.
+Figure 5b shows that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance.
+For example, at 5 Mbps, TX2 processing in its top three performance modes outperforms offloading; however, as TX2’s performance decreases beyond those modes, offloading becomes more advantageous.
+Across all scenarios, our models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.
 Key takeaway. At high network bandwidths, edge offloading becomes favorable as its computational advantage outweighs the reduced network delays.
-4.5. Impact of Request Rate
-Figure 5. Latency comparison between on-device processing and offloading to A2 for MobileNetV2 with varying request rates under 10 Mbps and 20 Mbps networks.
-Next, we demonstrate how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload.
-In this experiment, both the TX2 and Orin Nano are set to their maximum performance modes.
-We evaluate edge offloading under 10 Mbps and 20 Mbps network bandwidth.
-Figure 5 plots the end-to-end latency of each strategy as a function of request rate (RPS).
-When comparing Orin Nano with edge offloading to A2, we observe two distinct behaviors depending on the network bandwidth.
-At 10 Mbps, on-device processing using Orin Nano outperforms edge offloading across all request rates.
-However, under a 20 Mbps network, there exists a performance crossover point: On-device processing on Orin Nano provides lower latency under low load (RPS ¡ 60), but edge offloading becomes the faster option once the RPS exceeds 60.
-On the other hand, TX2 consistently exhibits the highest latency across all request rates due to its limited processing capability.
-Overall, the predicted end-to-end latencies of our models match closely with the observed latency, proving the accuracy of our models.
-This result shows that while both the network queueing delays (for offloading) and local processing queueing delay (for on-device processing) grow with $\lambda$, the rate at which delays increase differs depending on the hardware and network configurations.
-Under 10 Mbps, the network queueing delay increases at a faster rate than the device queueing delay as $\lambda$ increases, favoring on-device processing.
-Conversely, under 20 Mbps, device queuing delays dominate at higher $\lambda$, shifting the advantage to offloading.
-These competing scaling effects validate the observation we discussed in Lemma 3.1.
-Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency under faster networks.
-4.6. Impact of Multi-Tenancy
-Figure 6. Latency comparison across execution strategies for InceptionV4 under a 100 Mbps network with varying number of co-located applications. Each application has RPS = 2.
-We now experimentally validate our model’s ability to predict average response time in multi-tenant settings using Orin Nano as the device and A2 as the edge server.
+IV-F Impact of Collaborative Processing
+To demonstrate our models’ ability to capture collaborative (i.e., split) processing behavior, we consider an example using MobilNetV2 and use the Orin Nano as the device and A2 as the edge under a 50 Mbps network.
+We evaluate three split points (SPs) where the model is progressively offloaded to the edge and compare them to full on-device execution.
+For each split point, the initial portion of the neural network is executed locally on the device, while the resulting intermediate output and remaining computation are offloaded to the edge.
+Figure 6a plots the observed latencies of each strategy alongside model predictions.
+To compute the end-to-end latency, we use the combined analytic models for device and edge processing as discussed in Section III-E.
+As shown, the latency increases with the amount of compute offloaded.
+This is because the intermediate results of later layers grow in size, adding to the network transfer overhead.
+As seen, for MobileNetV2 and the chosen network configuration, local processing is faster than split processing.
+Across all split points, the model-predicted latencies closely align with observed values.
+Note that other models will exhibit other types of behaviors with split processing, which we omit due to space constraints.
+Key takeaway. Our models can be extended to capture both compute and network delays in device-edge collaborative processing.
+IV-G Impact of Request Rate
+Figure 6b shows how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload under 10 Mbps and 20 Mbps network bandwidth, with both the TX2 and Orin Nano set to their maximum performance modes.
+We observe two distinct behaviors depending on network bandwidth.
+At 10 Mbps, on-device processing using Orin Nano outperforms offloading across all RPS.
+However, at 20 Mbps, a performance crossover occurs: on-device processing is faster at low load (RPS $\leq$ 60), while offloading becomes more efficient beyond that point.
+In contrast, processing on TX2 consistently exhibits higher latency across all RPS due to its limited performance.
+In all cases, as the system approaches saturation at high RPS, queue lengths show increased variability from the Poisson arrival process, leading to greater fluctuations in queuing delays and response times.
+Overall, the predicted end-to-end latencies of our models match closely with the observed latency within confidence intervals, demonstrating the accuracy of our models.
+This result validates the competing scaling effects of $\lambda$ discussed in Lemma III.1.
+Although both the network queuing delays (for offloading) and local processing queuing delay (for on-device processing) grow with $\lambda$, the rate at which the delays increase differs depending on the hardware and network configurations.
+At 10 Mbps, the network queuing delay grows faster than the device queuing delay as $\lambda$ increases, favoring on-device processing.
+Conversely, at 20 Mbps, device queuing delays become dominant at higher $\lambda$, shifting the advantage toward offloading.
+Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency on faster networks.
+IV-H Impact of Multi-Tenancy
+We now experimentally validate the accuracy of our models in multi-tenant settings, using Orin Nano as the device and A2 as the edge server.
 We vary the number of applications deployed on the edge, each serving InceptionV4 requests, with all models loaded in memory concurrently.
-Each application receives 2 requests per second from its corresponding client.
-We report the average offloading latency across all clients and compare it with on-device processing.
-Figure 6 compares the end-to-end latency of offloading with on-device processing as the number of co-located applications increases.
-When only a single application is deployed on the edge, offloading yields lower latency than on-device processing.
-As the edge server hosts more applications, the latency of offloading increases due to resource contention across applications.
-We observe a performance crossover point at 18 co-located applications, beyond which on-device processing becomes more efficient.
-Note that the on-device processing time remains constant regardless of the number of concurrent applications as devices operate independently.
+Each application receives 2 RPS from its corresponding client.
+We report the average offloading latency and compare it with on-device processing.
+Figure 6c compares the average latency of offloading with on-device processing as the number of co-located applications increases.
+We observe that when only a single application is deployed on the edge, offloading yields lower latency than on-device processing.
+However, as the number of co-located applications increases, offloading latency also rises.
+This increase is caused by resource contention among applications because the A2 GPU does not support isolation.
+As a result, there is a crossover point at 18 applications, beyond which on-device processing becomes more efficient.
+Note that the on-device processing latency remains stable regardless of the number of concurrent applications as each device operates independently.
 In all cases, our models closely predict the latency for both execution strategies.
-Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on the edge server.
-5. Models in Action
-In this section, we present a resource manager that leverages our analytic models to adaptively decide between offloading and on-device processing.
-We outline its decision algorithm and demonstrate its effectiveness through two case studies that showcase how the resource manager adapts to network variability and workload interference in multi-tenant edge environments.
-5.1. Model-Driven Adaptive Resource Management
-We implement a resource manager that utilizes our analytic models to adaptively switch between on-device and edge processing.
-The manager operates locally on the device and periodically collects runtime metrics such as network bandwidth, edge server load, and request arrival rate.
-For each epoch, it uses the models to estimate the average latency of both execution strategies and chooses the option with the lower predicted average latency to execute requests.
-Algorithm 1 outlines the decision-making process for selecting the execution strategy in the resource manager.
-At Lines 1-2, it predicts on-device average execution latency based on the current arrival rate and device processing speed.
-At Lines 3-6, it computes the predicted average latency for offloading to each edge server, considering the server load, service rates, and the device’s network conditions.
-Finally, at Lines 7-11, it executes the request using the best execution strategy based on these predictions.
+Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on edge servers lacking GPU isolation.
+V Models in Action
+In this section, we present a resource manager that leverages our analytic models to adaptively switch between offloading and on-device processing.
+We describe its decision algorithm and demonstrate its effectiveness through two case studies showcasing how the models enable adaptation to network variability and multi-tenant interference at the edge.
+V-A Model-Driven Adaptive Resource Management
+The resource manager runs locally on the device and periodically collects runtime metrics, including network bandwidth, edge server load, and request arrival rate.
+At each epoch, it inputs these metrics into our models to estimate the average latency of both execution strategies and selects the one with the lower predicted latency to execute requests.
+Algorithm 1 outlines the decision-making process for selecting the execution strategy.
+Lines 1-2 predict the on-device average execution latency based on the current arrival rate and the estimated service time on the device.
+Lines 3-6 compute the predicted average latency for offloading to each edge server, considering server load, service rates, and network conditions.
+Finally, Lines 7-11 execute the request using the strategy with the lower predicted latency.
 Input:
@@
 // Predict edge offloading latency
-3 $T_{net}^{req}\leftarrow M/M/1(\lambda_{dev},B/D_{req})+D_{req}/B_{dev}$
-4 $T_{net}^{res}\leftarrow M/M/1(\lambda_{edge,E},B/D_{res})+D_{res}/B_{edge}$
+3 $T_{net}^{req}\leftarrow M/M/1(\lambda_{dev},B/D_{req})+D_{req}/B$
+4 $T_{net}^{res}\leftarrow M/M/1(\lambda_{edge,E},B/D_{res})+D_{res}/B$
 5 for $E\leftarrow 1$ to $\mathcal{E}$ do
@@
 10       $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
-       return offload($E^{*}$) ;
-        // Offload to server $E^{*}$
+       return offload($E^{*}$) // Offload to $E^{*}$
 11
-Algorithm 1 Model-Driven Adaptive Offloading Algorithm
-5.2. Case 1: Fluctuating Network Conditions
-Figure 7. Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
-We first demonstrate how the resource manager adapts to changes in network conditions.
-We use the TX2 as the device, A2 as the edge server, and MobileNetV2 as the workload, with requests arriving at a fixed rate of 10 RPS.
-Initially, the device has a network bandwidth of 20 Mbps.
-To simulate network variation, we reduce the bandwidth to 10 Mbps at t = 20, further reduce it to 2 Mbps at t = 40, and then increase it back to 20 Mbps at t = 60.
+Algorithm 1 Model-Driven Adaptive Offloading
+V-B Case 1: Fluctuating Network Conditions
+Figure 7: Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
+We first demonstrate how the resource manager leverages our models to adapt to changing network conditions.
+We use TX2 as the device, A2 as the edge, with MobileNetV2 requests arriving at a fixed rate of 10 RPS.
+Initially, the device has a network bandwidth of 20 Mbps to represent a stable 5G connection.
+To simulate real-world network variation, we reduce the bandwidth to 10 Mbps at t = 20 to reflect moderate signal degradation, further reduce it to 2 Mbps at t = 40 to emulate severe network congestion or poor coverage, and then restore it to 20 Mbps at t = 60 to represent signal recovery.
 Figure 7 shows the observed latency using the execution strategies selected by the resource manager, alongside the model-predicted latencies for both offloading and on-device processing.
@@
 Throughout this sequence, the models closely predict the observed values, validating their accuracy.
-Key takeaway. Our resource manager dynamically adapts to changing network conditions by switching between execution strategies.
-5.3. Case 2: Multi-tenant Edge Servers
-Figure 8.
-Observed latency under varying edge server workloads.
+Key takeaway. Our models enable the resource manager to dynamically adapt to changing network conditions by switching between execution strategies.
+V-C Case 2: Multi-tenant Edge Servers
+Figure 8: Observed latency under varying edge server workloads.
 At t = 80, the request rate of edge server 1 is increased from 10 RPS to 50 RPS.
 At t = 160, the request rate of edge server 2 is also increased to 50 RPS.
-Next, we demonstrate how our resource manager adapts to workload changes across multiple multi-tenant edge servers.
+Our second case study demonstrates how the resource manager leverages our models to adapt to workload changes across multiple multi-tenant edge servers.
 In this experiment, two edge servers, $E_{1}$ and $E_{2}$, each equipped with an A2 GPU, are hosting separate instances of a YOLOv8n application.
-Initially, the request rates are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while a TX2 running the resource manager has a request rate of $\lambda_{dev}=10$.
+Initially, the request rates of each server are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while the TX2 running the resource manager has a request rate of $\lambda_{dev}=10$.
 At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50.
@@
 In response, the resource manager adapts by offloading to the less loaded server $E_{2}$.
-At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and it shifts to on-device processing.
-This sequence demonstrates the resource manager’s ability to dynamically adapt to fluctuations in edge server workload.
-Key takeaway. Our resource manager dynamically adapts to workload changes in multi-tenant edge servers to maintain low latency.
-6. Related Work
+At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and the resource manager shifts to on-device processing.
+This sequence highlights how our model enables the resource manager to dynamically adapt to fluctuations in edge server workload.
+Key takeaway. Our models enable the resource manager to dynamically adapt to workload changes in multi-tenant edge servers to maintain low latency.
+VI Related Work
 Edge offloading.
-Edge offloading has become a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure (Davis et al., 2004; Jang et al., 2018; Lai et al., 2017; Li et al., 2024; Ogden et al., 2021; Ren et al., 2019; Mei et al., 2012; Liu et al., 2018; Chen et al., 2017; Wang et al., 2018; Zhang et al., 2024a; Tkachenko et al., 2023).
-For instance, Furion (Lai et al., 2017) offloads computation-intensive components of mobile AR pipelines to the edge.
-Jang et al.(Jang et al., 2018) propose an edge camera virtualization architecture to offload video analytics workloads.
-Li et al.(Li et al., 2024) introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints.
+Edge offloading is a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure [5, 8, 43, 44, 45, 6, 46, 47, 48, 49, 50, 51, 52].
+For example, Furion [43] offloads computation-intensive components of mobile AR pipelines to the edge.
+Li et al.
+[44] introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints.
 In contrast, our work uses a model-driven approach to reexamine whether offloading is advantageous in the era of accelerators.
-We extend our analysis to capture the combined effects of hardware characteristics, workload and network dynamics, and multi-tenancy in edge servers.
+Our models apply to a wide range of workloads and generalizes to device-edge collaborative processing as shown in the evaluation.
 Model-aware placement and resource allocation.
-Queueing models have been extensively used to predict the performance of traditional CPU-bound workloads (Jiang et al., 2021; Wang et al., 2024a; Tantawi and Towsley, 1985; Ambati et al., 2020; Gandhi and Suresh, 2019; Shen et al., 2011; Urgaonkar et al., 2005).
-Recently, there has been work focusing on modeling the performance of accelerators.
-For example, Reference (Merck et al., 2019) models the latency of GPU batching in DNN applications.
+Queuing models have been widely used to predict the performance of traditional CPU-bound workloads [53, 3, 54, 55, 56, 57, 20].
+More recently, there has been growing interest in modeling the performance of hardware accelerators.
+For example, Merck et al.
+[58] model the latency of GPU batching in DNN inference workloads.
 Liang et al.
-(Liang et al., 2023) model the performance of executing inference requests on various accelerators (e.g., TPU, GPU) and propose an online knapsack algorithm that uses the models for DNN model placement.
-In contrast, our models incorporate not only device hardware characteristics but also network delays and device-to-edge performance differences, enabling direct comparisons between on-device and edge offloading strategies.
-Collaborative inference.
-There is a growing body of work on collaborative inference, where the computation is split across multiple nodes in the network to reduce end-to-end latency (Almeida et al., 2022; Banitalebi-Dehkordi et al., 2021; Gao et al., 2022; Hu et al., 2019; Huang et al., 2020; Kang et al., 2017; Laskaridis et al., 2020; Li et al., 2018; Merck et al., 2019; Ng et al., 2024; Ogden and Guo, 2023; Tuli et al., 2023; Yang et al., 2022; Zhang et al., 2020, 2024b).
-Neurosurgeon (Kang et al., 2017) proposes regression models for layer-wise execution times prediction for DNN partitioning across network components.
-SPINN (Laskaridis et al., 2020) integrates early-exit networks with model partitioning across device and cloud, enabling graceful degradation under network instability.
-In this work, we focus on offloading entire inference requests and do not consider partitioned execution.
-7. Conclusions
-In this paper, we presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators.
+[29] develop performance models for inference on accelerators (e.g., GPU, TPU) and use them to guide DNN placement via an online knapsack algorithm.
+In contrast, our models are applicable not only to inference workloads but also to other accelerator-driven applications as shown in Section IV-D.
+They jointly capture device hardware characteristics, network delays, and device-to-edge performance differences, enabling direct comparison between on-device processing and edge offloading.
+Modeling distributed edge workloads.
+Several studies have developed analytic models for distributed device–edge–cloud environments.
+Raei et al.
+[59] analyze the impact of wireless signal strength on request rejection probability and mean response delay.
+FogQN [60] uses multi-class queuing networks to model fog and cloud computing to determine the optimal fraction of data processing executed at the cloud versus at fog servers.
+Loghin et al.
+[61] model MapReduce applications in hybrid edge–cloud environments.
+However, these works abstract away accelerator-specific queuing behavior and do not consider the effects of multi-tenancy at the edge.
+Our work complements existing efforts by addressing these missing dimensions through a unified queuing-based modeling framework that informs resource management decisions.
+VII Conclusions
+This paper presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators.
 We validated the accuracy of our model across a range of hardware, workloads, and network configurations.
-We developed a resource manager that leverages these predictions to dynamically adapt to changing network conditions and multi-tenant edge server workloads.
-As future work, we plan to extend our models to incorporate collaborative processing, i.e., performing a portion of request computation locally before offloading.
+We developed a resource manager to demonstrate how model predictions can be used to dynamically adapt the execution strategy in response to changing network conditions and multi-tenant edge server workloads.
+As future work, we plan to extend the framework to model continuous batching in LLM inference and to support multi-stage pipelines or DAGs with dependencies across stages.

### v3

@@
-To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing
-Abstract
+To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing In the Era of Accelerators
+Abstract.
 Computational offloading is a promising approach for overcoming resource constraints on client devices by moving some or all of an application’s computations to remote servers.
-With the advent of specialized hardware accelerators, client devices are now able to perform fast local processing of specific tasks, such as machine learning inference, reducing the need for offloading computations.
+With the advent of specialized hardware accelerators, client devices can now perform fast local processing of specific tasks, such as machine learning inference, reducing the need for offloading computations.
 However, edge servers with accelerators also offer faster processing for offloaded tasks than was previously possible.
-In this paper, we present an analytic and experimental comparison of on-device processing and edge offloading for a range of accelerator, network, and application workload scenarios, with the goal of understanding when to use local on-device processing and when to offload computations.
-We present models that leverage analytical queuing results to capture the effects of dynamic factors such as the performance gap between the device and edge server, network variability, server load, and multi-tenancy on the edge server.
-We experimentally demonstrate the accuracy of our models for a range of hardware and application scenarios and show that our models achieve a mean absolute percentage error of 2.2% compared to observed latencies.
-We use our models to develop an adaptive resource manager for intelligent offloading and show its efficacy in the presence of variable network conditions and dynamic multi-tenant edge settings.
-publicationid: pubid:
-I Introduction
+In this paper, we present an analytic and experimental comparison of on-device processing and edge offloading for a range of accelerator, network, multi-tenant, and application workload scenarios, with the goal of understanding when to use local on-device processing and when to offload computations.
+We present models that leverage analytical queuing results to derive explainable closed-form equations for the expected end-to-end latencies of both strategies, which yield precise, quantitative performance crossover predictions that guide adaptive offloading.
+We experimentally validate our models across a range of scenarios and show that they achieve a mean absolute percentage error of 2.2% compared to observed latencies.
+We further use our models to develop a resource manager for adaptive offloading and show its effectiveness under variable network conditions and dynamic multi-tenant edge settings.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
+1. Introduction
 Over the past decade, cloud computing has become the predominant approach for running online services in domains ranging from finance to entertainment.
 In recent years, a new class of online services has emerged that requires low-latency processing to meet end-user requirements.
-Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference [1].
-Edge computing, which is a complementary approach to cloud computing, is a promising method to meet the needs of such latency-sensitive applications.
+Examples of such applications include autonomous vehicles, interactive augmented and virtual reality (AR/VR), and human-in-the-loop machine learning inference (Mosqueira2022).
+Edge computing, a complementary approach to cloud computing, is a promising method for meeting the needs of such latency-sensitive applications.
 Conventional wisdom has held that since edge resources are deployed at the edge of the network and are closer to end-users, edge-native processing can provide lower end-to-end latencies than cloud processing.
-However, recent research [2, 3] has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud.
-In particular, since edge clusters are resource-constrained relative to the cloud, research has shown resource bottlenecks can arise in the presence of time-varying workloads and increase the processing latency at the edge, negating its network latency advantage.
-Further, such bottlenecks can arise even at moderate levels of utilization, requiring careful resource management to maintain the benefits of edge-native processing [3].
+However, recent research (Ali2021; WangB2024) has shown that this conventional wisdom does not always hold—in certain scenarios, edge processing can yield worse end-to-end latency despite its network latency advantage over the cloud.
+In particular, since edge clusters are resource-constrained relative to the cloud, research has shown that resource bottlenecks can arise under time-varying workloads and increase the processing latency at the edge, negating its network latency advantage and requiring careful resource management to maintain its benefits (WangB2024).
 A related trend is the emergence of computational offloading from local devices to remote resources such as those at the edge.
 [^1] Computational offloading to the cloud is also common, but we focus on edge offloading in this paper.
-Computational offloading [4, 5] involves using nearby edge resources to take on some, or all, of an application’s processing demands.
-Such offloading is performed to overcome resource constraints on a local device by leveraging more abundant processing capacity at edge servers.
-Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers [6].
-In this case, local devices pay a so-called mobility penalty [7], which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead.
-Offloading is also common for Internet-of-Things (IoT) devices that have limited on-board processing abilities and rely on remote processing to implement data processing tasks [8].
-In these offloading scenarios, conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
+Computational offloading (Satya2009; Davis2004) involves using nearby edge resources to take on some, or all, of an application’s processing demands.
+Such offloading is performed to overcome resource constraints on a device by leveraging the more abundant processing capacity at edge servers.
+Computational offloading is well-established in mobile computing where mobile devices, which are resource- or battery-constrained, offload local computations to edge servers (Ren2019).
+In this case, local devices pay a so-called mobility penalty (Satyanarayanan2019), which is the additional network latency to access edge servers, but then benefit from faster edge processing that outweighs this latency overhead.
+Conventional wisdom holds that remote edge processing is faster than local processing despite the network overhead of accessing remote resources due to significant resource constraints at local devices.
 With the advent of hardware accelerators in recent years, the conventional wisdom about the benefits of computational edge offloading for local devices needs to be rethought.
-For example, the rise of the so-called AI PC with onboard AI accelerators [9] has enabled local AI processing for many applications without a need for edge offloading.
-Similarly, mobile phones are equipped with neural accelerators [10] that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally [11].
+For example, the rise of the so-called AI PC with onboard AI accelerators (intel_ai_pc) has enabled local AI processing for many applications without the need for edge offloading.
+Similarly, mobile phones are equipped with neural accelerators (apple_neural_engine) that enable efficient local inference, while IoT devices such as cameras have specialized accelerators to perform visual tasks (e.g., facial recognition) locally (netatmo_camera).
 At the same time, edge servers can also be equipped with more powerful accelerators, such as GPUs, enabling offloaded tasks to also run more efficiently or faster than before.
@@
 (ii) Are there still scenarios where edge offloading has an advantage over local processing?
-(iii) What role does a fast (or slow) network between the device and the edge play in terms of whether to offload or not offload?
+(iii) How does multi-tenancy at edge servers affect the decision to offload or execute on-device?
 Motivated by the above questions, this paper takes an analytic model-driven approach to compare local and edge processing in the era of accelerators.
 Our hypothesis is that there is no a priori clear advantage of edge processing over on-device processing in the presence of accelerators, and that the choice of where to perform the processing depends on hardware, network, and workload characteristics.
-In developing our analytic models and validating this hypothesis, this paper makes the following contributions.
-  • We use queuing theory-based models to model the behavior of device and edge accelerators when processing requests for local on-device processing and edge offloading scenarios.
-Our modeling follows a two-level approach: it takes estimated service times obtained through profiling or other prediction techniques as input and embeds them into queuing models to obtain end-to-end response time predictions.
-Using these models, we develop closed-form analytic bounds to determine when one approach outperforms the other under different hardware, network, and workload scenarios.
-We also develop analytic models for comparing on-device processing to multi-tenant edge servers that serve multiple clients.
-Although our primary focus is on offloading, we also show how the models can be extended to account for device-edge collaborative (i.e., split) processing.
-  • We experimentally validate our analytic models and bounds using a variety of device and edge accelerators, network configurations, and workloads including deep neural networks (DNNs) and recurrent neural networks (RNNs).
-Our results show that our models are able to predict the performance behavior of on-device processing and edge offloading under various scenarios, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%.
-  • We develop a dynamic resource manager that incorporates our models and demonstrate how they can be leveraged by this resource manager to adaptively switch between on-device and edge processing under time-varying workloads and under dynamic multi-tenancy scenarios.
-Our results show that it effectively adapts to changing network conditions and fluctuating request rates at edge servers.
-II Background
+The novelty of this work lies in deriving closed-form equations for the expected end-to-end latency of each strategy, which are explainable and require no workload-specific training.
+These equations yield precise, quantitative performance-crossover predictions that can be embedded in OS schedulers and resource managers to dynamically decide whether to run an application component on the device or at the edge.
+In developing our analytic models and validating our hypothesis, this paper makes the following contributions.
+  • We develop queuing theory–based models to model the behavior of device and edge accelerators when processing requests in on-device and edge offloading scenarios.
+Our modeling adopts a two-level approach: it first uses estimated service times obtained through profiling or prediction techniques as input, and then embeds them into queuing models to predict end-to-end response times.
+We develop closed-form analytic bounds using these models to determine when one approach outperforms the other under different hardware, network, and workload scenarios.
+We further extend the analysis to model multi-tenant edge servers serving multiple clients, and demonstrate how the models can be extended to capture device–edge collaborative (i.e., split) processing.
+  • We experimentally validate our models and bounds across diverse device and edge accelerators, network configurations, and workloads spanning multiple model architectures, including deep neural network (DNN), recurrent neural network (RNN), and large language model (LLM).
+The results show that our models accurately predict on-device and offloading performance, achieving a 2.2% mean absolute percentage error, with 91.5% of predictions within $\pm$5% and all within $\pm$10% of the observed latency.
+  • We develop a resource manager that leverages our models to adapt between on-device processing and edge offloading in response to real-world dynamics.
+We present two case studies demonstrating that our models enable effective adaptation to changing network conditions and fluctuating request rates at multi-tenant edge servers.
+2. Background
 This section provides background on computational offloading, hardware accelerators, and on-device processing.
-TABLE I: Characteristics of common programmable accelerators.
-Accelerator Peak Power (W) Memory Cost Target Workload
-Intel Movidius Myriad X VPU 1 – 2.5 2.5 MB of SRAM \$59 Computer vision and CNN inference
-Google Edge TPU 2 8 MB of SRAM \$75 ML Inference with TensorFlow Lite
-NVIDIA Jetson TX2 15 8 GB of LPDDR4 \$399 AI and GPU workload
-NVIDIA Jetson Orin Nano 15 8 GB of LPDDR5 \$499 AI and GPU workload
-NVIDIA A2 GPU 60 16 GB of GDDR6 \$1399 AI and GPU workload
-II-A Cloud and Edge Offloading
+2.1. Computational Offloading
 Computational offloading is a well-known approach where client devices that are resource-constrained offload some or all of their application workload to a remote server.
-In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks [4, 5].
-While such edge offloading involves a network hop to the edge server (also known as the mobility penalty [7]), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than the devices.
-In the case of IoT devices or ML applications, offloading to the cloud is also common where the abundant processing capacity of the cloud platform is utilized for processing the application workload [12].
-For example, large language model (LLM) inference uses GPU clusters in the cloud to answer user queries, while IoT devices use cloud storage and processing to perform analytic and data processing tasks.
-In general, computational offloading offers benefits when local processing is constrained and the cost of incurring the additional network latency outweighs the benefits of accessing faster processing in the edge or the cloud.
-II-B Hardware Accelerators
-In recent years, a wide range of programmable accelerators has been developed to accelerate a range of compute tasks, particularly machine learning compute.
-Such accelerators have been deployed on client devices, edge, and cloud servers, enabling faster processing of accelerated tasks at each tier.
-Table 1 depicts the characteristics of several commonly deployed device and edge accelerators, ranging from low-end to high-end options.
-Specifically, the Intel Movidius Myriad X Vision Processing Unit (VPU) is designed to accelerate vision-based applications [13].
-Google Edge Tensor Processing Unit (TPU) optimizes ML inference with the TensorFlow Lite framework on the edge [14].
-NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision [15, 16, 17].
-For these specific workloads, programmable accelerators can provide substantial performance improvements, delivering server-like capabilities for some compute-intensive tasks even on low-end devices.
-With their small form factor and energy-efficient design, integrating accelerators into mobile devices has become increasingly common.
-For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device [10].
-With the presence of accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without the need to resort to edge offloading.
-Meanwhile, edge servers can also be equipped with GPUs, which serve as general-purpose accelerators for a range of tasks, or special-purpose accelerators for specific tasks.
+In the case of mobile devices such as AR/VR headsets, tablets, or smartphones, offloading can use nearby edge resources to perform latency-sensitive tasks (Satya2009; Davis2004).
+While such edge offloading involves a network hop to the edge server (also known as the mobility penalty (Satyanarayanan2019)), subsequent processing at the edge is assumed to be much faster since edge resources are much less constrained than those of devices.
+In general, computational offloading offers benefits when local processing is constrained and the benefits of accessing faster processing at the edge or in the cloud outweigh the cost of additional network latency.
+2.2. Hardware Accelerators
+In recent years, a wide variety of accelerators have been developed to speed up diverse compute tasks, particularly machine learning workloads.
+Table 1 depicts the characteristics of several commonly deployed device and edge accelerators.
+For example, NVIDIA TX2, Orin Nano, and A2 are all tailored for AI workloads such as deep learning and computer vision.
+With their small form factor and energy-efficient design, accelerators are increasingly integrated into mobile devices.
+For example, Apple devices include a Neural accelerator to accelerate machine learning tasks locally on the device (apple_neural_engine).
+With accelerators, fast on-device processing of tasks such as machine learning inference has become feasible without resorting to edge offloading.
+Meanwhile, edge servers can also be equipped with even more powerful GPUs.
 This enables offloaded tasks to also be accelerated at the edge, making edge offloading useful for more compute-intensive tasks that a device may not be able to handle using local resources.
-II-C On-device Processing versus Edge Offloading
-As noted above, the availability of accelerators on client devices has enabled on-device processing of many tasks without the need to offload those tasks to the edge or to the cloud.
-Since local processing does not incur any network hops, it will often be faster than remote processing which incurs an additional network latency.
+Table 1. Characteristics of common hardware accelerators.
+Accelerator Power Memory Workload
+Google Edge TPU (edgetpu2019) 2 W 8 MB ML Inference
+NVIDIA Jetson TX2 (tx22017) 15 W 8 GB GPU compute
+NVIDIA Jetson Orin Nano (orin2023) 15 W 8 GB GPU compute
+NVIDIA A2 GPU (a22021) 60 W 16 GB GPU compute
+2.3. On-device Processing versus Edge Offloading
+As noted above, the availability of accelerators on client devices has enabled on-device processing of many tasks without the need to offload those tasks to the edge.
+Since local processing does not incur any network hops, it will often be faster than remote processing which incurs additional network latency.
 However, since edge servers can also be equipped with GPUs and other accelerators, in some compute-intensive scenarios, edge processing may still provide benefits over local processing on device accelerators.
-Thus, in the era of accelerators, local on-device processing may be faster in some scenarios, while offloading to edge accelerators may still be faster in other cases (after accounting for the network latency).
-Even with extensive offline profiling of execution times on both the device and the edge, selecting the optimal execution strategy remains challenging due to dynamic real-world conditions.
-On the one hand, mobile devices are subject to network variability due to fluctuating wireless signal strength, interference, and congestion in shared communication channels.
-Since edge offloading requires transmitting input data to an edge server, network fluctuations introduce unpredictability in network delays, affecting its overall performance.
-Additionally, mobile devices running applications with prolonged operating durations, such as live object detection on captured video, may experience battery constraints.
-When operating in low-power mode, the system may underclock its processing units to conserve energy, resulting in extended execution times.
-On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can result in interference across users and higher latencies.
-Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically based on the number and nature of tasks being offloaded at any given time.
-Compared to traditional cloud data centers, edge servers are often more resource-constrained, limiting their ability to elastically scale applications under resource pressure.
-Workload spikes in multi-tenant scenarios can increase queuing delays, significantly increasing end-to-end execution times.
-Moreover, edge servers are commonly multiplexed across different applications to maximize edge resources utilization.
-At the same time, GPUs and other programmable accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs [18], leading to performance interference between applications.
-As a result, execution times can become variable, making it challenging to determine whether to offload under dynamic workloads.
+Thus, in the era of accelerators, local on-device processing may be faster in some scenarios, while offloading to edge accelerators may still be faster in others (after accounting for network latency).
+Even with extensive offline profiling, selecting the optimal execution strategy remains challenging due to real-world dynamics.
+On the one hand, mobile devices are subject to network variability such as fluctuations in wireless signal strength.
+Since edge offloading involves transmitting input data, network fluctuations can introduce unpredictability in network delays, thereby affecting offloading performance.
+Additionally, devices running applications with prolonged operating durations such as live object detection may face battery constraints.
+In low-power mode, processors may be underclocked by the OS to conserve energy, extending on-device execution times.
+On the other hand, edge servers are multi-tenant systems shared by multiple devices, which can lead to interference among clients and higher latencies.
+Since devices may independently decide when and whether to offload tasks, the edge server’s load can fluctuate dynamically with the number and type of incoming requests.
+Unlike traditional cloud data centers, edge servers are often more resource-constrained and cannot elastically scale all applications under high load.
+As a result, workload spikes in multi-tenant scenarios can increase queuing delays and significantly extend end-to-end execution times.
+Moreover, edge servers are often multiplexed across multiple applications to maximize resource utilization.
+GPUs and other accelerators designed for edge servers often lack isolation mechanisms such as GPU virtualization that are available in cloud-based GPUs (MIG_nvidia), leading to performance interference between applications.
+Consequently, execution times at the edge can vary, making it challenging to decide whether to offload under dynamic workloads.
 The above observations motivate the primary research question addressed in this paper: Considering the hardware processing capabilities of the device and edge server, along with the workload characteristics and real-world dynamics, under what scenarios does one strategy outperform the other?
@@
 How do workload dynamics, network dynamics, and edge multi-tenancy impact each strategy and their relative performance with respect to one another?
-To address this question, the next section develops analytic models to capture the end-to-end execution time for each on-device and edge processing, capturing the key factors that impact application performance.
-III Model-driven Performance Comparison
-In this section, we develop analytical queuing models for edge offloading and on-device processing, and derive closed-form expressions for their expected latencies to compare the two approaches and identify when one outperforms the other.
-III-A Analytic Queuing Models
+To address these questions, the next section develops analytic models for the end-to-end execution time of edge offloading and on-device processing, capturing the key factors that impact application performance.
+3. Model-driven Performance Comparison
+In this section, we develop analytic queuing models for edge offloading and on-device processing, and derive closed-form expressions for their expected latencies to compare the two approaches and identify when one outperforms the other.
+3.1. Analytic Queuing Models
 (a) Edge offloading queuing model
 (b) On-device processing queuing model
-Figure 1: Modeling request execution using (a) edge offloading and (b) on-device processing.
-Our analytical models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems.
-Queuing theory has previously been used to model the performance of edge and cloud applications [2, 19, 20], and will serve as the foundation for analyzing edge offloading and on-device processing.
-As discussed below, we model the end-to-end latency of a request—the duration of time from when the task enters the first queue until it exits the system—for both edge offloading and on-device processing using queuing theory.
+Figure 1. Modeling request execution using (a) edge offloading and (b) on-device processing.
+Our analytic models are derived from queuing theory, a well-established mathematical tool for modeling response times experienced by requests in computing systems.
+Queuing theory has previously been used to model the performance of edge and cloud applications (Ali2021; Harchol2013; Urgaonkar2005), and here will serve as the foundation for analyzing the end-to-end latency of requests—from the time a task enters the system until it completes—under both edge offloading and on-device processing.
 Edge offloading.
-To model the end-to-end request execution time using edge offloading, we need to model the times spent by the request both at the device and at the edge server.
-We do so by modeling the device and the edge as two separate queuing systems as shown in Figure 1a.
-We assume that the request arrives at the device and is then forwarded (“offloaded”) to the edge server over the network.
-We need to model any network queuing delays and the network transmission time of the request from the device to the server.
-Upon arrival at the edge server, the request is queued for processing and is subsequently scheduled for execution.
-Since accelerators employ data parallelism to process requests, we model the accelerator at the edge server as a parallel system that processes each request with a degree of parallelism $k_{edge}$.
-Once the request processing completes, the result is sent back to the device, which incurs network queuing delay and transmission latency on the reverse path.
+To model the end-to-end latency of a request under edge offloading, we need to model the times spent by the request both at the device and at the edge server, which we model as two separate queuing systems as shown in Figure 1a.
+We assume that a request arrives at the device and is then forwarded (“offloaded”) to the edge over the network, incurring network queuing and transmission delays.
+Upon arrival at the edge, the request is queued for processing and is subsequently scheduled for execution.
+Since accelerators can process multiple requests in parallel, we model the edge accelerator as a parallel system with a degree of parallelism $k_{edge}$.
+After processing, the result is sent back to the device, which incurs network queuing and transmission delays on the reverse path.
 On-device processing.
 Similar to edge offloading, we model on-device request processing as a queuing system shown in Figure 1b.
-Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores of the device where they are processed with a degree of parallelism $k_{dev}$.
-Upon completion of processing, the result is returned to the local application and the request exits the system.
-This model captures an application where a client device generates tasks, utilizes its accelerator cores for local execution, and sends the result to the application.
+Incoming requests generated by the device first enter a local device queue and are then scheduled onto one of the accelerator cores, where they are processed with a degree of parallelism $k_{dev}$.
+When processing completes, the result is returned to the local application, and the request exits the system.
+This model captures an application where a client device generates tasks, uses its accelerator cores for local execution, and sends the results back to the application.
 In queuing theory, the processing time of a request at each queue that it encounters consists of two key components: queuing delay (i.e., the time spent waiting in the queue) and service time (i.e., the time taken to execute the request on the accelerator).
 Let $w_{dev}^{proc}$, $w_{edge}^{proc}$, $s_{dev}$, and $s_{edge}$ represent the wait times and request service times at the device and edge, respectively.
-Edge offloading additionally incurs network latency to transmit the request to the edge and the response back to the device, while on-device processing does not incur any network latency since all processing is local.
+Edge offloading additionally incurs network latency as requests are sent to the edge and results are returned to the device.
 Let $w_{dev}^{net}$ and $w_{edge}^{net}$ denote the respective network queuing delays on the device and the edge.
-Also, let $n_{req}$ denote the network delay of sending a request from the device to the edge, and $n_{res}$ be the network delay of sending a result from the edge to the device.
+Also, let $n_{req}$ denote the network transmission delay of sending a request from the device to the edge, and $n_{res}$ denote the transmission delay of the response along the reverse path.
 With the above notations, we now model the end-to-end latency of requests executed using each strategy:
-$T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (1)
+(1) $T_{edge}=w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
 and
-$T_{dev}=w_{dev}^{proc}+s_{dev}$ (2)
+(2) $T_{dev}=w_{dev}^{proc}+s_{dev}$
 where $T_{edge}$ and $T_{dev}$ denote the end-to-end request latency under edge offloading and on-device processing, respectively.
-III-B Deriving Service Times on Accelerators
-Our analytic models require the service time of a workload as input to compute end-to-end response times.
-The service time reflects the computational demand on the target accelerator and indicates how quickly a request is processed on a given platform.
-It can be obtained through empirical profiling or a deep learning model that predicts workload execution times on accelerators.
-For empirical profiling, many applications already support function-level performance logging or provide profiling hooks [21], which can be used to derive the service time of a request by analyzing the corresponding function’s execution metrics.
-Alternatively, the workload can be profiled on the target hardware using benchmarking tools [22, 23].
-In case of model-based prediction, a neural network can be trained to predict the service time of a model on a given target hardware.
-For example, prior work [24, 25, 26] estimates DNN inference latency by analyzing factors such as network structure and parameter size.
-For recurrent architectures such as Gated Recurrent Unit (GRU) networks and Long Short-Term Memory (LSTM) networks, existing latency estimation techniques [27, 28] incorporate factors such as temporal dependencies, hidden state dimensionality, and sequence unrolling behavior to accurately predict inference latency.
-While any of these prior techniques can be used with our analytic models, our experiments use a simple neural network model to predict the service times.
-III-C Deriving and Analyzing End-to-End Latency
+3.2. Deriving Service Times on Accelerators
+Our models require the service time of a workload as input to compute end-to-end response times.
+The service time reflects the computational demand on the target accelerator and represents the time needed to process a request on that platform.
+It can be obtained either through empirical profiling or via a predictive deep learning model.
+For profiling, many applications already provide function-level performance logging or profiling hooks (nvidia2023fastertransformer), which can be used to derive service times by analyzing execution metrics.
+Alternatively, the workload can be benchmarked directly on the target hardware using tools such as (Reddi2020; nvidiasmi).
+For model-based prediction, a neural network can be trained to predict the service time of a model on a given hardware.
+For example, prior work predicts DNN inference latency by analyzing network structure and parameter size (Wang2024; Yao2018; Kang2017), RNN inference latency by examining temporal dependencies and sequence unrolling behavior (Ji2024; Li23), and LLM inference latency by using internal layer embeddings to predict remaining output length (Rana2024).
+While any of these prior techniques can be used with our analytic models, in our experiments we adopt a simple neural network from (Kang2017) to predict service times.
+3.3. Deriving and Analyzing End-to-End Latency
 In this section we compute the expected latency for edge offloading and on-device processing.
-Our analysis focuses on applications that require processing results on the device such as AR/VR applications.
-Note that our models can also be extended to capture the performance of applications that require results on the edge.
+Our analysis focuses on applications that require results on the device, such as AR/VR workloads.
+While the models assume edge offloading returns results to the device which incurs network delay, they can be generalized to applications where results are consumed on the edge (e.g., IoT sensing) by omitting this network delay.
 In the following, we leverage classical queuing theory to fundamentally understand the conditions under which one strategy outperforms the other.
-III-C1 Comparing Edge Offload and On-Device Execution
-We begin our analysis by examining the conditions under which edge offloading results in higher latency compared to on-device processing.
+3.3.1. Comparing Edge Offload and On-Device Execution
+We begin our analysis by examining the conditions under which edge offloading results in higher average latency compared to on-device processing.
 For workloads that benefit from hardware acceleration such as those utilizing GPUs, processing times can be significantly reduced compared to processing on general-purpose CPUs.
 In the following discussion, we refer to such workloads as accelerator-driven workloads.
-While our analysis primarily focuses on DNN inference, the results generalize to other AI workloads as will be discussed in Section III-E.
-Table II summarizes the notation used in our models.
-TABLE II: Notations used in analytic models.
+While our analysis primarily focuses on DNN inference, a representative component commonly found in AR/VR applications, the results generalize to other AI workloads as discussed in Sec 3.5.
+Table 2 summarizes the notation used in our models.
+Table 2. Notations used in analytic models.
 Notation Description
@@
 $\mu^{proc}_{dev},\mu^{proc}_{edge}$ Service rate of device/edge
-Lemma III.1.
-For accelerator-driven workloads, edge offloading incurs a higher average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
+Lemma 3.0.
+For accelerator-driven workloads, edge offloading incurs a higher average end-to-end latency than on-device processing when
+(3) $s_{dev}-s_{edge}<\frac{\lambda}{\mu_{dev}^{net}(\mu_{dev}^{net}-\lambda)}+\frac{\lambda}{\mu_{edge}^{net}(\mu_{edge}^{net}-\lambda)}+\frac{D_{req}+D_{res}}{B}+\frac{1}{2}\left(\frac{1}{k_{edge}\mu_{edge}^{proc}-\lambda}-\frac{1}{k_{edge}\mu_{edge}^{proc}}\right)-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
 Proof.
 Edge offloading will offer a worse latency than on-device processing when $T_{device}<T_{edge}$. That is
-$w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (3)
-Let $D_{req}$ and $D_{res}$ denote the task and result payload size, respectively. Further, let $B$ denote the bandwidth of the device-to-edge network path. The above inequality reduces to
-$s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$ (4)
-As shown in [29], DNN inference on programmable accelerators often exhibits deterministic service times (denoted as $D$), as the number of operations per request remains constant.
-Since typical accelerators have a small degree of parallelism $k$ [30], the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process.
-While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ [31, 32].
+(4) $w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
+Let $D_{req}$ and $D_{res}$ denote the request and result payload size, respectively. Further, let $B$ denote the bandwidth of the device-to-edge network path. The above inequality reduces to
+(5) $s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$
+As shown in (Liang2023), DNN inference on accelerators exhibits deterministic service times (denoted as $D$), as the number of operations per request remains constant.
+Since typical accelerators have a small degree of parallelism $k$ (Amert2017), the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process.
+While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ (Shah2018; Shah2020).
 This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$.
-Its expected wait time can be derived from the Pollaczek-Khinchin (P-K) formula [33, 19] and is given by
-$E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$ (5)
+Its expected wait time can be derived from the Pollaczek-Khinchin (P-K) formula (Gross2008; Harchol2013) and is given by
+(6) $E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$
 On the other hand, network interfaces with a single controller can be modeled as $M/M/1/FCFS$ systems with expected wait times given by
-$E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$ (6)
-Let $k_{dev}$ and $k_{edge}$ denote the degree of parallelism seen at the accelerator on the device and the edge respectively.
-Applying these two results to (4), we get {dmath} s_dev - s_edge ¡ 1μdevnet- λdevnet - 1μdevnet + 1μedgenet- λedgenet - 1μedgenet + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λedge - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
-Offloading is only required when a request arrives, which gives us $\lambda^{net}_{dev}=\lambda$.
+(7) $E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$
+Let $k_{dev}$ and $k_{edge}$ denote the degree of parallelism seen at the accelerator on the device and the edge, respectively. Applying these two results to (5), we get
+(8) $s_{dev}-s_{edge}<\frac{1}{\mu_{dev}^{net}-\lambda_{dev}^{net}}-\frac{1}{\mu_{dev}^{net}}+\frac{1}{\mu_{edge}^{net}-\lambda_{edge}^{net}}-\frac{1}{\mu_{edge}^{net}}+\frac{D_{req}+D_{res}}{B}+\frac{1}{2}\left(\frac{1}{k_{edge}\mu_{edge}^{proc}-\lambda_{edge}}-\frac{1}{k_{edge}\mu_{edge}^{proc}}\right)-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda_{dev}}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
+Offloading is required only when a request arrives, which gives us $\lambda^{net}_{dev}=\lambda$.
 For a queuing system to be stable, throughput (i.e., task completion rate) must equal the task arrival rate $\lambda$ under steady-state conditions.
 When tasks are offloaded to an edge server, results are transmitted to the device only upon task completion, which yields $\lambda_{edge}^{net}=\lambda$.
-The first two terms involving $\mu_{dev}^{net}$ on the RHS can be written as {dmath} 1μdevnet- λdevnet - 1μdevnet = λμdevnet(μdevnet- λ) whereas the second two terms involving $\lambda_{edge}^{net}$ can be written as {dmath} 1μedgenet- λedgenet - 1μedgenet = λμedgenet(μedgenet- λ).
 Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$.
-Substituting the above into (III-C1), we obtain {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof.
-∎
+Substituting the above into (8), we obtain
+(9) $s_{dev}-s_{edge}<\frac{\lambda}{\mu_{dev}^{net}(\mu_{dev}^{net}-\lambda)}+\frac{\lambda}{\mu_{edge}^{net}(\mu_{edge}^{net}-\lambda)}+\frac{D_{req}+D_{res}}{B}+\frac{1}{2}\left(\frac{1}{k_{edge}\mu_{edge}^{proc}-\lambda}-\frac{1}{k_{edge}\mu_{edge}^{proc}}\right)-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
+which completes the proof. ∎
 The above result implies the following two remarks.
-Remark III.1.
+Remark 3.2.
 On-device processing is likely to outperform edge offloading for workloads with lower computational demand.
 Explanation.
-Service time, which is the time spent processing a request on an accelerator, is determined by the computational demand of the workload divided by the processing capacity of the hardware.
-For any given hardware configuration, requests with lighter demand have to be processed much faster at the edge to overcome the network penalty for edge offloading to be faster than on-device processing.
-For requests with lower computational demands, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally due to the disparity in processing capacities between the device and the edge as the workload computational demand decreases.
-Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (III.1) will also become smaller since $\mu$ increases for lighter requests.
-The terms representing the network overheads remain unchanged since they depend on request size rather than the processing demand of the request.
-As a result, the inequality in (III.1) is more likely to hold true for requests with lighter processing demand.
+Service time, the time spent processing a request on an accelerator, is determined by the workload’s computational demand divided by the hardware’s processing capacity.
+For any given hardware configuration, requests with lower computational demands have to be processed much faster at the edge to overcome the network penalty for edge offloading to outperform on-device processing.
+For requests with lower demand, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally as the disparity in processing capacity between the device and the edge becomes less impactful.
+Further, since the service rate is the inverse of service time in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right-hand side of the inequality in (3) will also become smaller since $\mu$ increases for lighter requests.
+Meanwhile, the terms representing the network overheads remain unchanged, as they depend on the payload size rather than the request’s processing demand.
+As a result, the inequality in (3) is more likely to hold for requests with lighter processing demand.
 ∎
-Remark III.2.
+Remark 3.3.
 On-device processing is likely to outperform edge offloading on slower networks or when workloads have larger request or result payloads.
 Explanation.
-The term $\frac{D_{req}+D_{res}}{B}$ in (III.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality.
-In addition, longer transmission times reduce the service rates of the NIC at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the RHS.
-As a result, the inequality is more likely to hold for slower networks or tasks with higher network overheads, making on-device processing faster than edge offloading.
+The term $\frac{D_{req}+D_{res}}{B}$ in (3) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right-hand side of the inequality.
+In addition, longer transmission times reduce the NIC service rates at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the right-hand side.
+As a result, the inequality is more likely to hold under slower networks or for tasks with larger payloads, making on-device processing faster than edge offloading.
 ∎
 Practical takeaways.
-Lemma III.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective.
-The bound and the resulting inequality depends on several factors: (i) relative processing speeds of the device and edge, denoted by $s_{dev}$, $s_{edge}$ (and also the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) network speed, (iii) overall workload denoted by request rate $\lambda$, (iv) network transmission overheads.
-In general, for any specific workload, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading the workload; otherwise, on-device processing becomes preferable to offloading.
+Lemma 3.1 establishes a bound on the average service time difference between the device and the edge for edge offloading to be more effective.
+The bound and the resulting inequality depend on several factors: (i) relative processing speeds of the device and the edge $s_{dev}$ and $s_{edge}$ (and the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) overall workload denoted by request rate $\lambda$, (iii) network transmission overheads.
+In general, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading; otherwise, on-device processing becomes preferable to offloading.
 Moreover, the request rate $\lambda$ impacts offloading decisions in two competing ways.
-First, as $\lambda$ increases, both the edge and the device experience longer queuing delays, with the edge queue growing more slowly due to its faster processing speed.
-Consequently, the benefits of offloading improve under high load.
-However, under slow network conditions or for applications with large payloads, a high $\lambda$ increases the network queuing delay (first two terms of the right side of the inequality) and may lead to network congestion, making on-device processing preferable.
-Thus, the optimal strategy depends on the workload’s payload size relative to both network bandwidth and the edge-device performance gap.
-Remark III.1 shows how the computational demand of the workload influences the execution strategy.
-For workloads with higher computational demand, edge offloading is more likely to outperform on-device processing, as the edge’s processing speed advantage becomes larger.
-Conversely, for workloads with lower computational demand, on-device processing is more efficient as the performance advantage of edge processing is likely outweighed by the associated network costs.
-Remark III.2 demonstrates how network delays affect the efficiency of edge offloading.
-With smaller network bandwidth or larger request/result payloads, both the network queuing delay and transmission time increase.
-As a result, edge offloading becomes less efficient, while on-device processing is unaffected since it does not rely on network resources.
+First, as $\lambda$ increases, both the edge and the device experience longer queuing delays, though the edge queue grows more slowly due to its faster processing speed, which favors offloading.
+Second, under slow networks or for tasks with large payloads, a high $\lambda$ increases network queuing delay, which penalizes offloading.
+Consequently, the optimal strategy depends on the workload’s payload size relative to network bandwidth and the performance gap between the device and the edge.
+Remark 3.2 indicates that edge offloading is likely to outperform on-device processing for workloads with high computational demand as the edge’s processing speed advantage dominates, while on-device processing is more efficient for lighter workloads since network overhead can outweigh the edge’s advantage.
+Remark 3.3 shows that edge offloading becomes less efficient with limited bandwidth or large payloads, which increase network queuing and transmission delays, whereas on-device processing remains unaffected.
 Extending to collaborative processing.
-Collaborative (i.e., split) processing involves partially processing requests locally on the client’s device before offloading them to a remote edge server to minimize the amount of data sent over the network [34, 35, 36, 12, 26, 37, 38].
+Collaborative (i.e., split) processing involves partially processing requests locally on the device before offloading them to a remote edge server to minimize the amount of data sent over the network (Almeida2022; Banitalebi2021; Hu2019; Huang2020; Kang2017; Li2018; Yang2022; Chen2025).
 To compute the end-to-end response time of such a strategy, we combine our analytic models for on-device processing and edge offloading into a tandem queuing network that accounts for both the partial local computation and the subsequent transmission and processing delays at the edge server.
@@
 Then, the intermediate output of size $D_{inter}$ passes through the components modeled in Figure 1a, with a network transmission time of $D_{inter}/B$ and a service time $s^{\prime}_{edge}$ representing the remaining computation time on the edge.
-This combined model can then be used to estimate performance and inform decision-making when selecting between local, edge, or collaborative processing strategies.
-III-D Impact of Multi-Tenancy
-Figure 2: $m$ devices offloading requests to an edge server.
-While the analysis in Lemma III.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work.
+This combined model can then be used to estimate performance and inform decision-making when selecting between local, edge, or collaborative processing.
+3.4. Impact of Multi-Tenancy
+While the analysis in Lemma 3.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work.
 Hence, we extend our model-driven comparison to multi-tenancy scenarios.
 Modeling edge application multiplexing.
-In the case where each edge server is dedicated to a single client device, the workload (i.e., the request arrival rate) at the edge and the device $\lambda^{proc}_{dev}$ and $\lambda^{proc}_{edge}$ are identical.
-However, when an edge server is multiplexed across $m$ devices, its workload will be the sum of the individual device workloads under the offloading case, as shown in Figure 2.
+In the case where each edge server is dedicated to a single client device, the workload (i.e., the request arrival rate) at the edge and at the device $\lambda^{proc}_{dev}$ and $\lambda^{proc}_{edge}$ are identical.
+However, when an edge server is multiplexed across $m$ devices, its workload will be the sum of the individual device workloads under the offloading case, as shown below:
 This higher workload will increase the queuing delays at the edge and may cause the end-to-end offloading latency to exceed the on-device processing latency.
-Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process [39], it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload.
-That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the $i-th$ device.
+Since the request arrivals at each client device are governed by independent Poisson processes, and the composition of independent Poisson processes results in a new Poisson process (pishro2014), it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ equal to the sum of the workloads of all devices.
+That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the ithi-th device.
 We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary.
-For this case, we model each system as an $M/G/1$ system, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge.
+In this case, we model the edge system as an $M/G/1$ system, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the edge.
 This extension captures the potentially large variability in service times across applications in multi-tenant environments.
-If $s_{i}$ denotes the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times.
-That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ and the the the effective service rate of the edge $\mu_{edge}=1/s_{edge}$.
-Then the aggregate utilization of the edge server $\rho_{edge}$ is $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$.
+Let $s_{i}$ denote the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times.
+That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ with the effective service rate of the edge given by $\mu_{edge}^{proc}=1/s_{edge}$.
+The aggregate utilization of the edge server is then $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$.
 This leads us to the following lemma for the multi-tenant case.
-Lemma III.2.
-For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgekedgeμedgeprocVar[sedge]2(kedgeμedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
+Lemma 3.0.
+For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when
+(10) $s_{dev}-s_{edge}<\frac{\lambda_{dev}}{\mu_{dev}^{net}(\mu_{dev}^{net}-\lambda_{dev})}+\frac{\lambda_{edge}}{\mu_{edge}^{net}(\mu_{edge}^{net}-\lambda_{edge})}+\frac{D_{req}+D_{res}}{B}+\frac{\rho_{edge}+\lambda_{edge}k_{edge}\mu_{edge}^{proc}\text{Var}[s_{edge}]}{2(k_{edge}\mu_{edge}^{proc}-\lambda_{edge})}-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda_{dev}}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
 Proof.
-The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula [33, 19] {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (III-D) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into (4) completes the proof.
-∎
+The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by P-K formula (Gross2008; Harchol2013)
+(11) $E[w_{M/G/1}]=\frac{\rho+\lambda\mu\text{Var}[s]}{2(\mu-\lambda)}$
+Substituting (11) for $w_{edge}^{proc}$, (6) for $w_{dev}^{proc}$, and (7) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into (5) completes the proof. ∎
 Practical takeaways.
-The expected wait time in $M/G/1/FCFS$ systems is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays.
+The expected wait time in $M/G/1/FCFS$ systems is heavily influenced by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays.
 It is therefore important for application designers to consider service time variability when assigning applications to edge servers.
@@
 If such features were available, each device could be assigned to an isolated virtual GPU on the edge.
-In this special case, the offloading situation is reduced to a situation similar to that considered in Sec III-C1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device.
+In this special case, the offloading situation reduces to the one similar to that considered in Sec 3.3.1 where each virtual GPU provides $\frac{1}{m}$-th of the capacity to each device.
 However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
-III-E Discussion
+3.5. Discussion
 Generalizing to other accelerator-driven workloads.
-While the above analysis focuses on DNN workloads with deterministic service times modeled using an $M/D/1$ system, our modeling framework can be extended to other workloads with different service time characteristics by incorporating appropriate queuing results.
-For example, workloads with recurrent architectures such as GRUs and LSTMs often process variable-length inputs with service times that vary with the input sequence length.
-Similarly, early-exit DNNs allow requests to exit at different layers based on confidence thresholds, which results in variable service times.
-To capture such variability, our framework can adopt an $M/M/1$ queuing system where service times follow an exponential distribution with the expected wait time shown in (6).
-By selecting the appropriate queuing model, our framework can adapt to diverse workloads and provide accurate response time estimates as will be shown in Section IV-D.
-Currently, our framework does not account for workloads that violate the FCFS assumption, such as Large Language Model (LLM) inference with continuous batching where requests are dynamically batched, and extending the models to capture such behavior is a subject of future work.
+While the above analysis focuses on DNN workloads modeled with deterministic service times, our models can be extended to other accelerator-driven workloads with different service time characteristics by incorporating appropriate queuing formulations.
+For example, recurrent models have variable execution times that depend on input length.
+Similarly, LLM inference exhibits variable service times due to variations in the number of autoregressive decoding steps.
+To capture such variability, our models can adopt an $M/M/1$ system in which service times follow an exponential distribution, with the expected wait time given in (7).
+By selecting the appropriate queuing model, our models can adapt to diverse workloads and provide accurate response time estimates as shown in Sec 4.4.
 Model limitations.
@@
 Although no closed-form solution exists for the expected wait time in this setting, Marshall et al.
-[40] provide an upper bound {dmath} E[w_G/G/1] ≤λ(σa2+ σs2)2(1-ρ) where $\sigma_{a}$ and $\sigma_{s}$ are the variances for the interarrival and service distributions respectively.
-Additionally, the queuing results used in our analysis assume that each hardware queue has infinite capacity.
-For scenarios with bounded queue sizes, we can incorporate results for finite-buffer queuing systems [19] to model the impact of such capacity constraints.
-IV Model Validation
-In this section, we empirically evaluate how the key factors modeled in our analytic framework affect the performance of the two execution approaches and demonstrate that our models can accurately predict the response times across a wide range of configurations.
-IV-A Experimental Setup
+(marshall1968) provide an upper bound
+(12) $E[w_{G/G/1}]\leq\frac{\lambda(\sigma_{a}^{2}+\sigma_{s}^{2})}{2(1-\rho)}$
+where $\sigma_{a}^{2}$ and $\sigma_{s}^{2}$ are the variances of the request interarrival time and service time distributions, respectively.
+Additionally, the queuing results used in our analysis assume that each queue has infinite capacity.
+For scenarios with bounded queue sizes, we can incorporate results for finite-buffer queuing systems (Harchol2013) to model the impact of such capacity constraints.
+4. Model Validation
+In this section, we empirically validate the accuracy of our analytic models for predicting the end-to-end response times of edge offloading, on-device processing, and collaborative processing across a wide range of configurations.
+4.1. Experimental Setup
 Workloads.
-We validate our model’s ability to capture the impact of workload characteristics using a diverse set of workloads, including DNN-based models as well as recurrent models.
-For DNN workloads, we focus on cognitive assistant applications involving image classification and object detection.
-Specifically, we use three widely deployed DNN architectures, namely MobileNet, Inception, and YOLO, each with distinct computational demands and input sizes as shown below:
+We validate our model’s ability to capture the impact of workload characteristics using DNN, RNN, and LLM workloads.
+For DNN workloads, we focus on inference tasks for AR applications, specifically image classification, using three widely deployed architectures: MobileNet, Inception, and YOLO, each with distinct compute demands and input sizes:
 Model Input Shape Params (M) FLOPs (G)
@@
 YOLOv8n $[640,640,3]$ 3.2 8.7
-All inference requests are executed using TensorFlow 2.13.0 across the evaluated hardware.
-We use frames extracted from a 720p dash cam video as input to the models, with each frame processed either locally on the device or offloaded to the edge.
-Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process, which runs on a separate machine to avoid performance interference with the device during evaluation.
-For recurrent models, we focus on personal voice assistant applications involving sentiment classification of conversational utterances.
-We validate our models using two types of recurrent neural networks: a GRU model with lower computational demand and a more compute-intensive LSTM model, with their architectural specifications summarized below:
-Model Embedding Dim # Layers $\times$ Units Activation
-GRU 512 2 $\times$ 512 Sigmoid
-LSTM 1024 3 $\times$ 1024 Sigmoid
-Both models are implemented using TensorFlow’s Keras API.
-To simulate real-world user inputs, we use the Amazon Alexa Topical Chat dataset [41], a knowledge-grounded human-to-human conversation dataset consisting of spoken-style dialogues with an average utterance length of 19 tokens.
+All requests are executed using TensorFlow 2.13.0 on frames extracted from a 720p video.
+For RNN, we use a Long Short-Term Memory (LSTM) model implemented using TensorFlow’s Keras API for sentiment classification, with the Amazon Alexa Topical Chat dataset (Topical2023) as input.
+For LLM, we use the Llama-3.2-1B model (llama3_2_1B) run via llama.cpp (llama_cpp), with prompts from the Hugging Face ShareGPT dataset (sharegpt2024) as input.
+The specifications of both models are summarized below:
+Model Architecture Parameters # of Layers
+LSTM Recurrent 40 M 3
+Llama-3.2-1B Transformer 1.2 B 16
+Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that generates requests following a Poisson process, running on a separate machine to avoid performance interference with the device.
 Hardware.
-We use a diverse set of hardware platforms to reflect real-world deployment scenarios.
-For client devices with accelerators, we use the NVIDIA Jetson TX2 and Jetson Orin Nano, two commonly deployed platforms in IoT applications, with the TX2 being an earlier model and the Orin Nano offering better performance.
-For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU, a 40-core Intel Xeon E5-2660v3 CPU, and 256GB RAM.
-We also use a server with an NVIDIA RTX 4070 GPU, 8-Core AMD Ryzen 7 5700X CPU, and 32GB RAM.
+For client devices, we use the NVIDIA Jetson TX2 and Jetson Orin Nano, two commonly deployed platforms in IoT applications, with the former being an earlier model and the latter offering higher performance.
+For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU and 256GB RAM.
+We also use a server with an NVIDIA RTX 4070 GPU and 32GB RAM.
 Both edge servers are connected to a 1 Gbps network.
 These platforms span a wide performance range, from 1.3 TFLOPS (TX2) to 2.6 TFLOPS (Orin Nano), 4.5 TFLOPS (A2), and up to 29.1 TFLOPS (RTX 4070) in peak FP16 throughput.
+All platforms run Ubuntu 18.04.
+We use the Linux Traffic Control (TC) subsystem to emulate different network bandwidths.
 For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable.
-To estimate $k$, we empirically measure how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior.
-All platforms run Ubuntu 18.04, and we use the Linux Traffic Control (TC) subsystem to emulate different network bandwidths.
-IV-B Parameter Estimation
-Our model requires several parameters to compute the end-to-end response times.
-We now describe how these parameters are obtained in our evaluation and outline other practical methods for estimating them.
+We estimate $k$ by empirically measuring how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior.
+4.2. Parameter Estimation
+Our models require several parameters to compute the end-to-end response times.
+We now describe how they are obtained in our evaluation.
+Additional practical approaches for estimating them are discussed in Sec 3.2.
 Estimating average service times.
-Our system estimates workload service times through empirical profiling and a neural network–based prediction approach.
-We estimate service times for DNN workloads by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool [23], which reports per-process execution times on the accelerator.
-For GRU and LSTM workloads, we profile request latencies on a representative input set distinct from the test set and use the average latency from this profiling as the service time of the model.
-This profiling data is used for our edge-offloading and on-device processing experiments.
-For collaborative processing, a neural network is trained using profiling data to predict service time based on prior work discussed in Section III-B.
-Such a neural network approach avoids the extensive profiling overhead of all possible split configurations by predicting partial execution service times.
+We estimate service times through profiling and a neural network–based prediction approach.
+For DNN workloads, we monitor inference durations using the NVIDIA System Management Interface (nvidia-smi) tool (nvidiasmi), which reports per-process execution times on the accelerator.
+For RNN and LLM workloads, we profile request latencies on a representative input set separate from the test set and use the average latency as input to the models.
+This profiling data informs our edge-offloading and on-device processing experiments.
+For collaborative processing, a neural network is trained on profiling data to predict service time as discussed in Sec 3.2 to avoid the extensive profiling overhead of all possible split configurations by predicting partial service times.
 Estimating average network delays.
-Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf [42].
+We estimate transmission delay using the available network bandwidth, measured continuously in real time with iperf (iperfwebsite).
 In addition to queuing and transmission delays captured by our models, network delays also include propagation and processing delays.
@@
 The service rate $\mu$ can be estimated from system logs based on the number of completed requests within a given time interval.
-IV-C Impact of Workload Characteristics
-Our first experiment investigates how workload characteristics influence offloading decisions using three DNN workloads with distinct profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload).
+4.3. Impact of Workload Characteristics
+We first study how workload characteristics influence offloading decisions using three DNN models: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload).
 We compare on-device processing on TX2 and Orin Nano against offloading to an A2 GPU using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps.
-For each device, we test all available GPU frequency settings to emulate different device conditions.
-Figure 3: Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2; Right: Orin Nano vs A2.
-Figure 3 compares the average latency of on-device processing and edge offloading across workloads, devices, and GPU frequency levels.
-Figures 3a and 3b show that for MobileNetV2, local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels.
-The trend reverses for InceptionV4 as shown in Figures 3c and 3d, where offloading provides lower average latency across all device configurations.
-Lastly, Figures 3e and 3f shows the results for YOLOv8n.
-In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes despite YOLOv8n’s high compute demand, while Orin Nano achieves lower average latency than offloading across all performance modes.
-We observe similar results when offloading to RTX4070 and omit the details here due to space constraints.
-In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%, proving the accuracy of our models.
-Collectively, these empirical results align with the qualitative observations stated in Remark III.1 and Remark III.2.
-While the performance ratio between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands.
-For workloads with low computational demand (MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delays of offloading.
-However, for compute-intensive workloads (InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network delays.
-Moreover, workloads with large message payloads (YOLOv8n) favor on-device processing because it does not involve any network components and thus avoids any network delay.
+Figure 2. Latency comparison for DNN workloads: (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n.
+Figure 2 compares the average latency of on-device processing and edge offloading across workloads and devices.
+For each device, we test all available GPU frequency settings to emulate varying device conditions, with each data point representing one setting.
+Figures 2a and 2b show that for MobileNetV2, TX2 achieves lower average latency in its top three frequency levels, while Orin Nano consistently outperforms offloading across all frequency levels.
+The trend reverses for InceptionV4 as shown in Figures 2c and 2d, where offloading provides lower average latency across all device configurations.
+Lastly, Figures 2e and 2f show that despite YOLOv8n’s high compute demand, TX2 outperforms offloading in its top six frequency levels, while Orin Nano achieves lower average latency across all settings.
+We observe similar results when offloading to an RTX4070 GPU and omit the details here due to space constraints.
+Collectively, these results align with the qualitative observations stated in Remark 3.2 and Remark 3.3.
+While the performance ratio between device and edge is fixed, the absolute advantage of offloading scales with the workload’s computational demands.
+For MobileNetV2 (low compute), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delays of offloading.
+However, for InceptionV4 (high compute), the same performance ratio yields larger absolute improvements that outweigh the network delays.
+Moreover, YOLOv8n (large payloads) favors on-device processing because it avoids any network delay.
+In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5% of the observed latency and 100% within $\pm$10%, proving their accuracy.
 Key takeaway.
-Offloading provides lower latency when the edge’s absolute performance gain exceeds the associated network delay.
+Offloading provides lower average latency when the edge’s absolute performance gain exceeds the associated network delay.
 Our models accurately predict the latency for both strategies with a mean absolute percentage error of 2.2%.
-IV-D Impact of Complex Deep Learning Models
-Figure 4: Latency comparison across execution strategies for (a) GRU model and (b) LSTM model at 3 RPS under 10 Mbps.
-Figure 5: Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
-Figure 6: Latency comparisons across (a) different collaborative processing configurations, (b) request rates under 10 Mbps and 20 Mbps networks, and (c) numbers of co-located applications.
-In addition to the DNN workloads, our model also accurately captures the latency behavior of more complex deep learning workloads.
-Figure 4 compares the average latency of on-device processing on the Orin Nano against offloading to A2 for two recurrent models: GRU and LSTM, under a moderate load of 3 RPS.
-Since service times vary with input length in these models, we incorporate the $M/M/1$ queuing model into our analysis to account for this variability in latency prediction.
-Figure 4a shows that even for the lightweight GRU model, offloading offers lower latency due to the small input utterance length (on average less than 100 bytes), which incurs minimal network delay.
-Figure 4b shows that as the model complexity increases, the absolute performance gain of offloading increases, leading to a greater latency reduction.
+4.4. Impact of Complex Deep Learning Models
+Figure 3. Latency comparison of execution strategies for (a) LSTM and (b) Llama-3.2-1B models.
+In addition to DNN workloads, our model accurately captures the latency behavior of other deep learning workloads.
+Figure 3 compares the average latency of on-device processing on the Orin Nano against offloading to the A2 GPU for RNN and LLM workloads under moderate loads.
+Since service times vary across requests in these models, we use the $M/M/1$ queuing model to account for this variability.
+Figure 3a shows that for the LSTM model, offloading achieves lower average latency as the input utterances are short and their payloads incur only minimal network delay.
+In the LLM case shown in Figure 3b, which has a higher computational demand than LSTM, the latency reduction from offloading becomes even more pronounced, reaching up to 30 seconds.
 In both cases, our model predictions closely match the observed latencies.
-Key takeaway. Our models generalize to complex deep learning workloads beyond DNN inference by integrating appropriate queuing models.
-IV-E Impact of Network Conditions
-Figure 5 shows how network speed affects the effectiveness of each execution strategy.
-Figure 5a compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds.
-For both devices, we observe a performance crossover point where offloading becomes faster than on-device processing.
-The location of the crossover depends on both network bandwidth and the performance gap between the device and the edge.
+Key takeaway. Our models generalize to other deep learning workloads by integrating appropriate queuing models.
+4.5. Impact of Network Conditions
+Figure 4. Latency comparison under varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
+Figure 4a compares on-device processing on TX2 and Orin Nano with offloading to the RTX4070 GPU under varying network bandwidths.
+It shows that the performance crossover point depends on both bandwidth and the performance gap between the device and the edge.
 For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial.
-Figure 5b shows that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance.
-For example, at 5 Mbps, TX2 processing in its top three performance modes outperforms offloading; however, as TX2’s performance decreases beyond those modes, offloading becomes more advantageous.
-Across all scenarios, our models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.
-Key takeaway. At high network bandwidths, edge offloading becomes favorable as its computational advantage outweighs the reduced network delays.
-IV-F Impact of Collaborative Processing
-To demonstrate our models’ ability to capture collaborative (i.e., split) processing behavior, we consider an example using MobilNetV2 and use the Orin Nano as the device and A2 as the edge under a 50 Mbps network.
+Figure 4b shows that even for a single device, multiple crossover points may exist depending on bandwidth and the device’s performance.
+At 5 Mbps, TX2 in its top three frequency levels outperforms offloading, but as its performance drops beyond those levels, offloading becomes advantageous.
+In all cases, our models closely predict the crossover points, showing their accuracy in capturing the tradeoff between the computation speedup from edge offloading and network delay.
+Key takeaway. At high network bandwidths, edge offloading becomes favorable as its computational advantage outweighs the smaller network delays.
+4.6. Impact of Collaborative Processing
+Figure 5. Latency comparisons of execution strategies under different: (a) collaborative processing configurations, (b) request rates, and (c) numbers of co-located applications.
+To validate our models’ ability to capture collaborative (i.e., split) processing behavior, we split the execution of MobilNetV2 across Orin Nano and A2 over a 50 Mbps network.
 We evaluate three split points (SPs) where the model is progressively offloaded to the edge and compare them to full on-device execution.
-For each split point, the initial portion of the neural network is executed locally on the device, while the resulting intermediate output and remaining computation are offloaded to the edge.
-Figure 6a plots the observed latencies of each strategy alongside model predictions.
-To compute the end-to-end latency, we use the combined analytic models for device and edge processing as discussed in Section III-E.
+For each split point, the early neural network layers are executed on the device, and the intermediate output and remaining computation are offloaded to the edge.
+To compute the end-to-end latency, we use the combined analytic model discussed in Sec 3.5.
+Figure 5a plots the observed latencies of each configuration alongside model predictions.
 As shown, the latency increases with the amount of compute offloaded.
@@
 As seen, for MobileNetV2 and the chosen network configuration, local processing is faster than split processing.
+Note that other models will exhibit other types of behaviors with collaborative processing, which we omit due to space constraints.
 Across all split points, the model-predicted latencies closely align with observed values.
-Note that other models will exhibit other types of behaviors with split processing, which we omit due to space constraints.
-Key takeaway. Our models can be extended to capture both compute and network delays in device-edge collaborative processing.
-IV-G Impact of Request Rate
-Figure 6b shows how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload under 10 Mbps and 20 Mbps network bandwidth, with both the TX2 and Orin Nano set to their maximum performance modes.
-We observe two distinct behaviors depending on network bandwidth.
-At 10 Mbps, on-device processing using Orin Nano outperforms offloading across all RPS.
+Key takeaway. Our models can be extended to accurately predict the performance of device-edge collaborative processing.
+4.7. Impact of Request Rate
+Figure 5b shows how the request rate $\lambda$ affects offloading decisions for the MobileNetV2 workload under 10 Mbps and 20 Mbps networks, illustrating two distinct behaviors depending on the network bandwidth.
+At 10 Mbps, Orin Nano outperforms offloading across all RPS.
 However, at 20 Mbps, a performance crossover occurs: on-device processing is faster at low load (RPS $\leq$ 60), while offloading becomes more efficient beyond that point.
-In contrast, processing on TX2 consistently exhibits higher latency across all RPS due to its limited performance.
-In all cases, as the system approaches saturation at high RPS, queue lengths show increased variability from the Poisson arrival process, leading to greater fluctuations in queuing delays and response times.
-Overall, the predicted end-to-end latencies of our models match closely with the observed latency within confidence intervals, demonstrating the accuracy of our models.
-This result validates the competing scaling effects of $\lambda$ discussed in Lemma III.1.
-Although both the network queuing delays (for offloading) and local processing queuing delay (for on-device processing) grow with $\lambda$, the rate at which the delays increase differs depending on the hardware and network configurations.
+This result validates the competing scaling effects of $\lambda$ discussed in Lemma 3.1.
+Although both the network queuing delays (for offloading) and processing queuing delays grow with $\lambda$, the rate at which the delays increase differs depending on the hardware, workload, and network configurations.
 At 10 Mbps, the network queuing delay grows faster than the device queuing delay as $\lambda$ increases, favoring on-device processing.
 Conversely, at 20 Mbps, device queuing delays become dominant at higher $\lambda$, shifting the advantage toward offloading.
-Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency on faster networks.
-IV-H Impact of Multi-Tenancy
-We now experimentally validate the accuracy of our models in multi-tenant settings, using Orin Nano as the device and A2 as the edge server.
-We vary the number of applications deployed on the edge, each serving InceptionV4 requests, with all models loaded in memory concurrently.
-Each application receives 2 RPS from its corresponding client.
-We report the average offloading latency and compare it with on-device processing.
-Figure 6c compares the average latency of offloading with on-device processing as the number of co-located applications increases.
-We observe that when only a single application is deployed on the edge, offloading yields lower latency than on-device processing.
-However, as the number of co-located applications increases, offloading latency also rises.
-This increase is caused by resource contention among applications because the A2 GPU does not support isolation.
+Overall, the predicted latencies of our models closely match the observed latencies within confidence intervals, demonstrating their accuracy.
+Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower average latency on faster networks.
+4.8. Impact of Multi-Tenancy
+Figure 5c compares the average latency of the two strategies as the number of co-located InceptionV4 applications on the edge increases, where each application receives 2 RPS from its corresponding client.
+The result shows that when only a single application is deployed on the edge, offloading yields lower average latency than on-device processing.
+However, as the number of co-located applications increases, offloading latency increases as well.
+This increase is caused by resource contention among applications because the A2 GPU lacks isolation mechanisms.
 As a result, there is a crossover point at 18 applications, beyond which on-device processing becomes more efficient.
-Note that the on-device processing latency remains stable regardless of the number of concurrent applications as each device operates independently.
-In all cases, our models closely predict the latency for both execution strategies.
-Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on edge servers lacking GPU isolation.
-V Models in Action
-In this section, we present a resource manager that leverages our analytic models to adaptively switch between offloading and on-device processing.
-We describe its decision algorithm and demonstrate its effectiveness through two case studies showcasing how the models enable adaptation to network variability and multi-tenant interference at the edge.
-V-A Model-Driven Adaptive Resource Management
-The resource manager runs locally on the device and periodically collects runtime metrics, including network bandwidth, edge server load, and request arrival rate.
-At each epoch, it inputs these metrics into our models to estimate the average latency of both execution strategies and selects the one with the lower predicted latency to execute requests.
-Algorithm 1 outlines the decision-making process for selecting the execution strategy.
-Lines 1-2 predict the on-device average execution latency based on the current arrival rate and the estimated service time on the device.
-Lines 3-6 compute the predicted average latency for offloading to each edge server, considering server load, service rates, and network conditions.
-Finally, Lines 7-11 execute the request using the strategy with the lower predicted latency.
+Note that the on-device processing latency remains stable regardless of the number of concurrent applications, as each device operates independently.
+In all cases, our models closely predict the latencies for both strategies.
+Key takeaway. Offloading average latency increases with the number of co-located applications due to resource contention on edge servers lacking GPU isolation.
+5. Models in Action
+In this section, we demonstrate how a resource manager can leverage our models to adaptively switch between offloading and on-device processing.
+We outline its algorithm and demonstrate its effectiveness through two case studies that showcase how the models enable adaptation to network variability and to multi-tenant edge servers.
+5.1. Model-Driven Adaptive Resource Management
 Input:
 $D_{req},D_{res}$: Request/result payload size
+0.1em
 $s_{dev}^{proc},s_{edge}^{proc}$: Service times of device and edge
 $\lambda_{dev}$: Device task arrival rate
+0.1em
 $B$: Network bandwidth of device
+0.1em
 $\mathcal{E}$: Number of edge servers
+0.1em
 $\{\mu_{edge,E}^{proc}\}_{E=1}^{\mathcal{E}}$: Edge server aggregated service rates
 $\{\lambda_{edge,E}\}_{E=1}^{\mathcal{E}}$: Edge server aggregated arrival rates
+0.2em
 Output:
 Execution strategy (on-device or edge server $E^{*}$)
-// Predict on-device processing latency
+1
+1ex// Predict on-device processing latency
 2 $\mu_{dev}^{proc}\leftarrow 1/s_{dev}^{proc}$
@@
 9  return on_device_processing()
-9else
+10else
 11  $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
   return offload($E^{*}$) // Offload to $E^{*}$
-11
+12
 Algorithm 1 Model-Driven Adaptive Offloading
-V-B Case 1: Fluctuating Network Conditions
-Figure 7: Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
-We first demonstrate how the resource manager leverages our models to adapt to changing network conditions.
-We use TX2 as the device, A2 as the edge, with MobileNetV2 requests arriving at a fixed rate of 10 RPS.
-Initially, the device has a network bandwidth of 20 Mbps to represent a stable 5G connection.
-To simulate real-world network variation, we reduce the bandwidth to 10 Mbps at t = 20 to reflect moderate signal degradation, further reduce it to 2 Mbps at t = 40 to emulate severe network congestion or poor coverage, and then restore it to 20 Mbps at t = 60 to represent signal recovery.
-Figure 7 shows the observed latency using the execution strategies selected by the resource manager, alongside the model-predicted latencies for both offloading and on-device processing.
-At t = 0, the predicted offloading latency is lower than on-device processing, therefore the manager chooses to offload requests.
-At t = 20, despite the reduced bandwidth from 20 to 10 Mbps, the predicted latency for offloading remains lower, and the strategy remains unchanged.
-However, at t = 40, the drop to 2 Mbps causes the predicted offloading latency (45 ms) to exceed the predicted on-device processing time (21 ms), and the resource manager to switch to on-device processing.
-Finally, at t = 60, as the network returns to 20 Mbps, the resource manager switches back to offloading.
-Throughout this sequence, the models closely predict the observed values, validating their accuracy.
-Key takeaway. Our models enable the resource manager to dynamically adapt to changing network conditions by switching between execution strategies.
-V-C Case 2: Multi-tenant Edge Servers
-Figure 8: Observed latency under varying edge server workloads.
-At t = 80, the request rate of edge server 1 is increased from 10 RPS to 50 RPS.
-At t = 160, the request rate of edge server 2 is also increased to 50 RPS.
-Our second case study demonstrates how the resource manager leverages our models to adapt to workload changes across multiple multi-tenant edge servers.
-In this experiment, two edge servers, $E_{1}$ and $E_{2}$, each equipped with an A2 GPU, are hosting separate instances of a YOLOv8n application.
-Initially, the request rates of each server are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while the TX2 running the resource manager has a request rate of $\lambda_{dev}=10$.
-At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50.
+The resource manager runs locally on the device and periodically collects runtime metrics including network bandwidth, edge server load, and request arrival rate as discussed in Sec 4.2.
+At each epoch, it inputs these metrics into our models to estimate the average latency for each strategy, and selects the strategy with the lowest predicted latency to execute requests.
+Algorithm 1 outlines the decision-making process of the resource manager.
+Lines 1-2 predict the average latency of on-device processing based on the current arrival rate and the device’s estimated service time.
+Lines 3-6 compute the predicted offloading latency for each edge server, considering server load, service rates, and network conditions.
+Finally, Lines 7-11 execute the request using the strategy with the lower predicted latency.
+5.2. Case 1: Fluctuating Network Conditions
+Figure 6. Latencies under varying network bandwidth as the resource manager selects between the two strategies.
+We first demonstrate how the resource manager can leverage our models to adapt to dynamic network conditions, using TX2 as the device, A2 as the edge, with MobileNetV2 requests arriving at a fixed rate of 10 RPS.
+Figure 6 shows the observed latency using the execution strategies chosen by the resource manager, alongside the model-predicted latency of each strategy.
+Initially, the device is configured with a 20 Mbps network to represent a stable 5G connection.
+Since the predicted offloading latency is lower than on-device processing, the manager chooses to offload requests.
+To emulate real-world network variation, we reduce the bandwidth to 10 Mbps at t = 20 to reflect moderate signal degradation.
+Here, despite the reduced bandwidth from 20 to 10 Mbps, the predicted latency for offloading remains lower, and the execution strategy remains unchanged.
+At t = 40, we further reduce the bandwidth to 2 Mbps to emulate poor coverage.
+Since the predicted offloading latency (45 ms) now exceeds the predicted on-device processing latency (21 ms), the resource manager switches to on-device processing.
+Finally, we restore the bandwidth to 20 Mbps at t = 60 to represent signal recovery, and the resource manager switches back to offloading.
+Throughout this sequence, our models closely predict the observed latencies, validating their accuracy.
+Key takeaway. Our models enable dynamic adaptation to changing network conditions by switching between execution strategies.
+5.3. Case 2: Multi-tenant Edge Servers
+Figure 7. Latencies under varying edge server workloads as the resource manager adapts.
+We now showcase how the resource manager can leverage our models to adapt to workload changes across multi-tenant edge servers.
+We use two edge servers, $E_{1}$ and $E_{2}$, each serving YOLOv8n inference requests using an A2 GPU.
+Initially, the server request rates are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while the TX2 running the resource manager has a request rate of $\lambda_{dev}=10$.
+Figure 7 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing.
+At t = 0, since the predicted latency for offloading to $E_{1}$ is the lowest, the resource manager offloads requests to $E_{1}$.
+At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50 to emulate workload fluctuations.
+In response, the resource manager adapts by offloading to the less loaded server $E_{2}$.
 Similarly, at t = 160, we increase $\lambda_{E_{2}}$ to 50.
-Figure 8 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing.
-Initially, from t = 0 to t = 80, the predicted latency for offloading to $E_{1}$ is the lowest, therefore the resource manager offloads requests to $E_{1}$.
-However, at t = 80, as the workload on $E_{1}$ increases, its predicted latency rises.
-In response, the resource manager adapts by offloading to the less loaded server $E_{2}$.
-At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and the resource manager shifts to on-device processing.
-This sequence highlights how our model enables the resource manager to dynamically adapt to fluctuations in edge server workload.
+Now, as the predicted latency of on-device processing becomes the lowest when both servers have a request rate of 50 RPS, the resource manager shifts to on-device processing.
+This sequence highlights how our model enables the resource manager to dynamically adapt to fluctuations in edge server workloads.
 Key takeaway. Our models enable the resource manager to dynamically adapt to workload changes in multi-tenant edge servers to maintain low latency.
-VI Related Work
+6. Related Work
 Edge offloading.
-Edge offloading is a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure [5, 8, 43, 44, 45, 6, 46, 47, 48, 49, 50, 51, 52].
-For example, Furion [43] offloads computation-intensive components of mobile AR pipelines to the edge.
+Edge offloading is widely adopted in many applications to enable resource-constrained devices to perform compute-intensive tasks by leveraging edge resources (Davis2004; Jang2018; Lai2017; Li2024; Ogden2021; Ren2019; Mei2012; Liu2018; zhou2017; Wang2018; Zhang2024MILCOM; Tkachenko2023; Ogden2023; Zilic2025).
+For example, Furion (Lai2017) selectively offloads compute-intensive components of mobile AR pipelines to the edge.
 Li et al.
-[44] introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints.
-In contrast, our work uses a model-driven approach to reexamine whether offloading is advantageous in the era of accelerators.
-Our models apply to a wide range of workloads and generalizes to device-edge collaborative processing as shown in the evaluation.
-Model-aware placement and resource allocation.
-Queuing models have been widely used to predict the performance of traditional CPU-bound workloads [53, 3, 54, 55, 56, 57, 20].
+(Li2024) introduce an evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints.
+In contrast, our work uses a model-driven approach to provide end-to-end latency predictions and quantitative performance crossover predictions for accelerator-driven workloads and generalizes to device-edge collaborative processing.
+Model-aware resource allocation.
+Queuing models have been widely used to predict the performance of traditional CPU-bound workloads (Jiang2021; WangB2024; Tantawi1985; Ambati2020; Gandhi2019; Shen2011; Urgaonkar2005).
 More recently, there has been growing interest in modeling the performance of hardware accelerators.
 For example, Merck et al.
-[58] model the latency of GPU batching in DNN inference workloads.
+(Merck2019) model GPU batching latency in DNN inference workloads.
 Liang et al.
-[29] develop performance models for inference on accelerators (e.g., GPU, TPU) and use them to guide DNN placement via an online knapsack algorithm.
-In contrast, our models are applicable not only to inference workloads but also to other accelerator-driven applications as shown in Section IV-D.
-They jointly capture device hardware characteristics, network delays, and device-to-edge performance differences, enabling direct comparison between on-device processing and edge offloading.
-Modeling distributed edge workloads.
+(Liang2023) develop performance models for inference on accelerators to guide DNN placement via an online knapsack algorithm.
+In contrast, our models apply not only to inference workloads but also to other accelerator-driven workloads as shown in Sec 4.4.
+Modeling distributed workloads.
 Several studies have developed analytic models for distributed device–edge–cloud environments.
-Raei et al.
-[59] analyze the impact of wireless signal strength on request rejection probability and mean response delay.
-FogQN [60] uses multi-class queuing networks to model fog and cloud computing to determine the optimal fraction of data processing executed at the cloud versus at fog servers.
+FogQN (Tadakamalla2018) uses multi-class queuing networks to model fog and cloud computing to determine the optimal fraction of data processing executed at the cloud versus at fog servers.
 Loghin et al.
-[61] model MapReduce applications in hybrid edge–cloud environments.
-However, these works abstract away accelerator-specific queuing behavior and do not consider the effects of multi-tenancy at the edge.
+(Loghin2019) model MapReduce applications in hybrid edge–cloud environments.
+However, they abstract away accelerator-specific queuing behavior and do not account for multi-tenancy at the edge.
 Our work complements existing efforts by addressing these missing dimensions through a unified queuing-based modeling framework that informs resource management decisions.
-VII Conclusions
-This paper presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators.
-We validated the accuracy of our model across a range of hardware, workloads, and network configurations.
-We developed a resource manager to demonstrate how model predictions can be used to dynamically adapt the execution strategy in response to changing network conditions and multi-tenant edge server workloads.
-As future work, we plan to extend the framework to model continuous batching in LLM inference and to support multi-stage pipelines or DAGs with dependencies across stages.
+7. Conclusions
+This paper presented analytic models for on-device processing and edge offloading with accelerators, providing performance predictions for adaptive execution decisions.
+We validated their accuracy across a range of scenarios and developed a resource manager that applies these predictions to adapt to real-world dynamics.
+As future work, we plan to extend the models to support multi-stage pipelines represented as DAG structures, capturing inter-stage dependencies.